### v1

# Combining Crown Structures for Vulnerability Measures

Abstract
Over the past decades, various metrics have emerged in graph theory to grasp the complex nature of network vulnerability. In this paper, we study two specific measures: (weighted) vertex integrity (wVI) and (weighted) component order connectivity (wCOC). These measures not only evaluate the number of vertices required to decompose a graph into fragments, but also take into account the size of the largest remaining component. The main focus of our paper is on kernelization algorithms tailored to both measures. We capitalize on the structural attributes inherent in different crown decompositions, strategically combining them to introduce novel kernelization algorithms that advance the current state of the field. In particular, we extend the scope of the balanced crown decomposition provided by Casel et al. [DBLP:conf/esa/Casel0INZ21] and expand the applicability of crown decomposition techniques.
In summary, we improve the vertex kernel of VI from $p^{3}$ to $p^{2}$, and of wVI from $p^{3}$ to $3(p^{2}+p^{1.5}p_{\ell})$, where $p_{\ell}<p$ represents the weight of the heaviest component after removing a solution. For wCOC we improve the vertex kernel from $\mathcal{O}(k^{2}W+kW^{2})$ to $3\mu(k+\sqrt{\mu}W)$, where $\mu=\max(k,W)$. We also give a combinatorial algorithm that provides a $2kW$ vertex kernel in FPT-runtime when parameterized by $r$, where $r\leq k$ is the size of a maximum $(W+1)$-packing. We further show that the algorithm computing the $2kW$ vertex kernel for COC can be transformed into a polynomial algorithm for two special cases, namely when $W=1$, which corresponds to the well-known vertex cover problem, and for claw-free graphs. In particular, we show a new way to obtain a $2k$ vertex kernel (or to obtain a 2-approximation) for the vertex cover problem by only using crown structures.
keywords:
Crown Decomposition, Kernelization, Vertex Integrity, Component Order Connectivity
category:
[…]
1 Introduction
In the study of graph theory different scales have emerged over the past decades to capture the complex nature of network vulnerability. While the main focus is on the connectivity of vertices and edges, subtle aspects of vulnerability such as the number of resulting components, the size distribution of the remaining components, and the disparity between them are becoming increasingly interesting. Our focus in this study is on two specific ways to measure vulnerability: (weighted) vertex integrity and (weighted) component order connectivity. These measures not only evaluate the number of vertices required to decompose a graph into fragments, but also take into account the size of the largest remaining component. Incorporating these aspects provides a more comprehensive understanding of network resilience.
Informally, vertex integrity (VI) models finding the right balance between removing vertices and keeping small connected parts of a graph. More formally, given a graph $G=(V,E)$ and a number $p\in\mathds{N}$, VI tries to find a set of vertices $S\subseteq V$ such that $|S|$ plus the size of the biggest component in the remaining graph is at most $p$. In the vertex weighted version (wVI), the goal is bounding the total weight of the removed vertices plus the weight of the heaviest component by $p$. This problem was introduced by Barefoot et al. [barefoot1987vulnerability] as a way to measure vulnerability of communication networks. Recently, it has drawn the interest of the parameterized complexity community due to its status as a natural parameter that renders numerous NP-hard problems amenable to fixed parameter tractability (FPT). This means that for these problems, solutions can be computed within a time frame represented by $f(p)\cdot n^{\mathcal{O}(1)}$, where $f$ is a computable function [DBLP:conf/isaac/LampisM21]. It is interesting to see how vertex integrity relates to other well-known measures of network structure. It imposes greater constraints compared to metrics such as treedepth, treewidth, or pathwidth, as the vertex integrity of a graph serves as an upper bound for these parameters. However, it encompasses a wider range of scenarios compared to vertex cover, where a vertex cover of a graph is an upper bound for its vertex integrity. This makes it a key in understanding how to efficiently solve problems in the world of network analysis.
The measure component order connectivity (COC) can be seen as the refined version of VI. Given a graph $G=(V,E)$ and two parameters $k,W\in\mathds{N}$, the goal of COC is to remove $k$ vertices such that each connected component in the resulting graph has at most $W$ vertices — also known in the literature as the $W$-separator problem or $\alpha$-balanced separator problem, where $\alpha\in(0,1)$ and $W=\alpha|V|$. In the vertex-weighted version (wCOC), the goal is to remove vertices of total weight at most $k$ such that the weight of the heaviest remaining component is at most $W$. An equivalent view of this problem is to search for the minimum number of vertices required to cover or hit every connected subgraph of size $W+1$. In particular, $W=1$ corresponds to covering all edges, showing that the COC is a natural generalization of the vertex cover problem.
The focus of the paper is on kernelization algorithms tailored for both weighted and unweighted versions of VI and COC when parameterized by $p$ and $k+W$, respectively. Kernelization algorithms can be thought of as formalized preprocessing techniques aimed to reduce optimization problems. Of particular interest in this work are crown decompositions, which are generally used as established structures for safe instance reduction - where "safe" means that there is no negative impact on the corresponding optimization problem. Essentially, a crown decomposition partitions the vertex set into distinct components: crown, head, and body. Here, the head acts as a separator between the crown and the body. This structural arrangement becomes useful when specific relationships between the head and the crown are required. Such relationships ultimately enable us to shrink instances by eliminating these designated parts from the graph. The properties of this structural layout, coupled with its existence guarantees depending on the instance size, enable the development of efficient kernelization algorithms for different problem domains. Notably, crown decompositions have also recently found utility in approximation algorithms for graph packing and partitioning problems [DBLP:conf/esa/Casel0INZ21]. For further exploration of crown decompositions, including their variations and applications, we recommend the comprehensive survey paper by Jacob et al. [DBLP:journals/algorithms/JacobMR23].
Our methods take advantage of the structural characteristics found in various crown decompositions, leading to the development of new kernelization algorithms that improve the state of the art. In essence, this work expands upon the applications of the balanced crown decomposition introduced by Casel et al. [DBLP:conf/esa/Casel0INZ21], specifically by integrating different crown decompositions into this framework.
Related Work:
Recently, the vertex integrity problem (VI) was extensively studied as a structural graph parameter [DBLP:conf/esa/BentertHK23, DBLP:journals/tcs/GimaHKKO22, DBLP:journals/algorithmica/GimaO24, DBLP:conf/isaac/LampisM21]. Gima et al. [DBLP:conf/walcom/GimaHKM0O24] conducted a systematic investigation into structural parameterizations computing the VI and wVI which was further extended by Hanaka et al. [DBLP:journals/corr/abs-2402-09971]. Additionally, there are notable results concerning special graph classes [DBLP:journals/dam/BaggaBGLP92, clark1987computational, DBLP:journals/algorithmica/DrangeDH16, DBLP:journals/dam/KratschKM97, DBLP:journals/ijcm/LiZZ08]. In our context, regarding related work on VI and wVI, Fellows and Stueckle presented an algorithm that solves the problem in $\mathcal{O}(p^{3p}n)$ [fellows1989immersion]. This was subsequently improved by Drange et al. [DBLP:journals/algorithmica/DrangeDH16], even for the weighted case, to $\mathcal{O}(p^{p+1}n)$. In the same paper, they presented the first vertex-kernel of size $p^{3}$ for both VI and wVI.
We turn now our attention to the related work of COC and wCOC. It is unlikely that kernelization algorithms can be achieved by considering $k$ or $W$ alone in polynomial time. Indeed, $W=1$ corresponds to a hard problem, which shows that $W$ (alone) is not a suitable parameter. For the parameter $k$, the problem is $W[1]$-hard even when restricted to split graphs [DBLP:journals/algorithmica/DrangeDH16]. These lower bounds lead to the study of parameterization by $k+W$. The best known algorithm with respect to these parameters finds a solution in time $n^{O(1)}\cdot 2^{\mathcal{O}(\log(W)\cdot k)}$ [DBLP:journals/algorithmica/DrangeDH16]. Unless the exponential time hypothesis fails, the authors prove that this running time is tight up to constant factors, i.e., there is no algorithm that solves the problem in time $n^{\mathcal{O}(1)}\cdot 2^{o(\log(W)\cdot\text{OPT})}$. The best known approximation algorithm has a multiplicative gap guarantee of $\mathcal{O}(\log(W))$ to the optimal solution with a running time of $n^{\mathcal{O}(1)}\cdot 2^{\mathcal{O}(W)}$ [DBLP:journals/corr/Lee16c]. They showed that the superpolynomial dependence on $W$ may be needed to achieve a polylogarithmic approximation. Using this algorithm as a subroutine, the vertex integrity can be approximated within a factor of $\mathcal{O}(\log(\text{OPT}))$, where OPT is the vertex integrity.
Regarding kernelization algorithms, there is a sequence of results which successively improve the vertex-kernel of COC. The first results came from Chen et al. [DBLP:journals/tcs/ChenFSWY19] and Drange et al. [DBLP:journals/algorithmica/DrangeDH16], who provided a $\mathcal{O}(kW^{3})$ and  $\mathcal{O}(k^{2}W+kW^{2})$ kernel, respectively. The result of Drange et al. also holds for wCOC and is the only result for this case. This was improved simultaneously by Xiao [DBLP:journals/jcss/Xiao17a] as well as by Kumar and Lokshtanov [DBLP:conf/iwpec/KumarL16] to a $\mathcal{O}(kW^{2})$ kernel. These works also provide the first $\mathcal{O}(kW)$ kernels, but with different constants and running times. Kumar and Lokshtanov [DBLP:conf/iwpec/KumarL16] present a $2kW$ kernel in a running time of $n^{\mathcal{O}(W)}$ by using linear programming (LP) methods with an exponential number of constraints (cf. LABEL:fig::LPWSep). The runtime can be improved to $2^{\mathcal{O}(W)}\cdot n^{\mathcal{O}(1)}$ as already mentioned in the book of Fomin et al. [fomin2019kernelization] (Section 6.4.2). Roughly speaking, the idea is to use the ellipsoid method with separation oracles to solve the linear program, where the separation oracle uses a method called color coding to find violated constraints that makes it tractable in $W$. Note that if $W$ is a constant this $2kW$ kernel is a polynomial time kernel improving on some previous results. This includes for instance the improvement of the $5k$ kernel provided by Xiao and Kou [DBLP:conf/tamc/XiaoK17] to a $4k$ kernel for the well-studied $P_{2}$-covering problem, where a $P_{2}$ is a path with 2 edges. The first linear kernel in both parameters in polynomial time, i.e. an $\mathcal{O}(kW)$ vertex kernel, is presented by Xiao [DBLP:journals/jcss/Xiao17a], who provide a $9kW$ vertex kernel. Finally, this was improved recently by Casel et al. [DBLP:conf/esa/Casel0INZ21] to a $3kW$ vertex kernel, which also holds for a more general setting. Namely, to find $k$ vertices in a vertex weighted graph such that after their removal each component weighs at most $W$. Note that the weights of the separator, i.e. the chosen $k$ vertices, play no role in this problem compared to wCOC. With the exception of the $\mathcal{O}(k^{2}W+kW^{2})$ vertex kernel of Drange et al. [DBLP:journals/algorithmica/DrangeDH16], all achieved vertex kernels essentially use crown structures.
Our Contribution:
The provided running times are based on an input graph $G=(V,E)$. We improve the vertex kernel for VI from $p^{3}$ to $3p^{2}$ in time $\mathcal{O}\left(\log(p)|V|^{4}|E|\right)$. For wVI, we improve it to $3(p^{2}+p^{1.5}p_{\ell})$ in time $\mathcal{O}\left(\log(p)|V|^{4}|E|\right)$, where $p_{\ell}<p$ represents the weight of the heaviest component after removing a solution.
To better explain the results of COC and wCOC, we introduce the problem of a maximum $\lambda$-packing for $\lambda\in\mathds{N}$. Given a (vertex weighted) graph $G$, this problem aims to maximize the number of disjoint connected subgraphs, each of size (or weight) at least $\lambda$. It is worth noting that the size of a maximum $(W+1)$-packing serves as a lower bound on $k$, since each element in the packing contains at least one vertex of a solution of COC (or wCOC) — in terms of linear programming, it is the dual of COC (cf. LABEL:fig::LPWSep).
For wCOC we improve the vertex kernel of $\mathcal{O}(k^{2}W+kW^{2})$ to $3\mu(k+\sqrt{\mu}W)$ in time $\mathcal{O}\left(r^{2}k|V|^{2}|E|\right)$, where $\mu=\max(k,W)$ and $r\leq|V|$ is the size of a maximum $(W+1)$-packing. For the unweighted version, we provide a $2kW$ vertex kernel in an FPT-runtime of $\mathcal{O}(r^{3}|V||E|\cdot r^{\min(3r,k)})$, where $r\leq k$ is the size of a maximum $(W+1)$-packing. In particular, for the vertex cover problem (where $W=1$), the kernelization result yields a $2k$ vertex kernel, thereby implying a $2$-approximation. That is, if we consider kernelization for COC within polynomial time to achieve a vertex kernel of size $ckW$ for some constant $c$, the optimal constant achievable is $c=2$, unless the unique games conjecture fails [KHOT2008335]. Comparing this result with the state of the art, disregarding the FPT-runtime aspect, we improve upon the best-known polynomial algorithm, achieving a kernel of size $3kW$ [DBLP:conf/esa/Casel0INZ21]. A $2kW$ vertex kernel is also presented in [DBLP:conf/iwpec/KumarL16], albeit with an exponential runtime using linear programming methods, which, as mentioned, can be enhanced to an FPT-runtime regarding parameter $W$. In contrast, our result is entirely combinatorial and has an FPT-runtime in the parameter of a maximum $(W+1)$-packing $r\leq k$. We further show that the algorithm computing the $2kW$ vertex kernel for COC can be transformed into a polynomial algorithm for two special cases. The first case arises when $W=1$, i.e., for the vertex cover problem. Here, we provide a new method for obtaining a vertex kernel of $2k$ (or obtaining a $2$-approximation) using only crown decompositions. The second special case is for the restriction of COC to claw-free graphs.
Regarding these special cases, until 2017, a $2k$ vertex kernel for $W=1$ was known through crown decompositions, albeit computed using both crown decompositions and linear programming. Previously, only a $3k$ vertex kernel was known using crown decompositions alone. In 2018, Li and Zhu [DBLP:journals/tcs/LiZ18] provided a $2k$ vertex kernel solely based on crown structures. They refined the classical crown decomposition, which possessed an additional property allowing the remaining vertices of the graph to be decomposed into a matching and odd cycles after exhaustively applying the corresponding reduction rule. In contrast, our algorithm identifies the reducible structures, which must exist if the size of the input graph exceeds $2k$. Unfortunately, we were unable to transform the FPT-runtime algorithm into a polynomial-time algorithm in general. Nevertheless, we believe that our insights into the structural properties of certain crown decompositions potentially pave the way to achieving this goal.
Lastly, all missing details, i.e. omitted proofs and algorithms as well as more detailed explanations can be found in the appendix.
2 Preliminaries
In this section, we briefly discuss terminology related to graphs and parameterized complexity that we use in the extended abstract (the complete terminology is given in LABEL:appendix::prelim). Additionally, we introduce the crown structures that are utilized throughout this paper.
Graph and Parameterized Terminology
We use a standard terminology for graphs $G=(V,E)$. The following is not necessarily common: For sets of vertex sets $\mathcal{V}\subseteq 2^{V}$ we abuse notation and use $V(\mathcal{V})=\bigcup_{Q\in\mathcal{V}}Q$ and $N(\mathcal{V}):=\left(\bigcup_{Q\in\mathcal{V}}N(U)\right)\setminus V(\mathcal{V})$. We define $\mathbb{CC}(G)\subseteq 2^{V}$ as the connected components of $G$ as vertex sets. Let $G=(V,E,w\colon V\to\mathds{N})$ be a vertex weighted graph. For $V^{\prime}\subseteq V$ and $G^{\prime}\subseteq G$ we define $w(V^{\prime})$ and $w(G^{\prime})$ as $\sum_{v\in V^{\prime}}w(v)$ and $\sum_{v\in V(G^{\prime})}w(v)$, respectively.
For parameterized complexity we use the standard terminology, which is also used, for example, in [downey2012parameterized, fomin2019kernelization]. Of particular interest in this work are kernelizations, which can be roughly described as formalized preprocessings. More formally, given an instance $(I,k)$ parameterized by $k$, a polynomial algorithm is called a kernelization if it maps any $(I,k)$ to an instance $(I^{\prime},k^{\prime})$ such that $(I^{\prime},k^{\prime})$ is a yes-instance if and only if $(I,k)$ is a yes-instance, $|I^{\prime}|\leq g(k)$, and $k^{\prime}\leq g^{\prime}(k)$ for computable functions $g,g^{\prime}$.
Crown Decompositions
We present a more general variant of crown decomposition that also captures commonly used crown compositions or expansions, which we explain in a moment. This novel decomposition can be easily derived from existing results of [DBLP:conf/esa/Casel0INZ21], but to the best of our knowledge has never been used in this form before. (See Figure 1 (left) for an illustration.)
Definition 2.1 (demanded balanced expansion and weighted crown decomposition).
For a graph $G=(A\cup B,E,w)$, a partition $A_{1}\cup A_{2}$ of $A$, an injective function $f\colon\mathbb{CC}(G[B])\to A$, demands $D=\{d_{a}\}_{a\in A}$ with $d_{a}\in\mathds{N}$ for each $a\in A$ and $y\in\mathds{N}$ the tuple $(A_{1},A_{2},y,f,D)$ is a demanded balanced expansion if
  1. 1.
    $w(Q)\leq y$ for each $Q\in\mathbb{CC}(G[B])$
  2. 2.
    $f(Q)\in N(Q)$ for each $Q\in\mathbb{CC}(G[B])$
  3. 3.
    $N(f^{-1}(A_{1}))\subseteq A_{1}$
  4. 4.
    $w(a)+w\left(f^{-1}(a)\right)\begin{cases}>d_{a}-y+1\text{ for each }a\in A_{1}\\ \leq d_{a}+y-1\text{ for each }a\in A_{2}\end{cases}$
If the demands are the same for each $a\in A$, e.g. $d_{a}=x$ for every $a\in A$ with $x\in\mathds{N}$, then we call $(A_{1},A_{2},y,f,x)$ a DBE. Furthermore, we call $(C,H,f)$ a ($x,y$) crown decomposition (($x,y$)-CD) if it is a $(H,\varnothing,y,f,x+y-1)$ DBE with $H=A$ and $C=f^{-1}(H)$.
[…]
$\mathfrak{R}$$H$$C$$\leq\lambda$$\lambda<$$\leq 3\lambda$
Figure 1:
Left: Let $A=\{a_{1},a_{2},a_{3},a_{4}\}$ be ordered in a top down manner and $w(Q)=1$ for every $Q\in\mathbb{CC}(G[B])$. Then, ($(\{a_{1},a_{2}\},\{a_{3},a_{4}\},1,f,\{3,2,4,4\})$ is a DBE, where the assignment $f$ are depicted with corresponding colored bold edges.
Right: $\lambda$-balanced crown decomposition [DBLP:conf/esa/Casel0INZ21]
In Definition 2.1, the vertex set $A_{1}$ typically represents the structure targeted for reduction and can be seen as the ’head’ of a crown decomposition. The components within $\mathbb{CC}(G[B])$ assigned to $A_{1}$, denoted by $f^{-1}(A_{1})$, form the ’crown’. In the context of optimization problems such as the vertex cover problem, a successful reduction often involves the head being part of an optimal solution. Subsequently, upon its removal, the crown — separated by the head — no longer needs to be considered.
The classical crown decomposition corresponds to a $(1,1)$-CD, the $q$-expansion for $q\in\mathds{N}$ to a $(q,1)$-CD, the weighted crown decomposition for $x\in\mathds{N}$ to a $(x,x)$-CD, and for $x,y\in\mathds{N}$ a $(A_{1},A_{2},y,f,x)$ DBE is called a balanced expansion. The essential new structural property of our DBE are the varying demands for $A$.
Let $G=(A\cup B,E)$ be a graph. For $A^{\prime}\subseteq A$ we define $\mathcal{B}_{A^{\prime}}$ as the components $Q\in\mathbb{CC}(G[B])$ with $N(Q)\subseteq A^{\prime}$. The following theorem gives the runtime to find a DBE and the existence guarantee of $A_{1}$ depending on the size or weight in the graph.
Theorem 2.2 (demanded balanced expansion).
Let $G=\left(A\cup B,E,w\right)$ be a graph with no isolated components in $\mathbb{CC}(G[B])$, $y\geq\max_{Q\in\mathbb{CC}(G[B])}w(Q)$ and $D=\{d_{a}\}_{a\in A}$ demands with $d_{a}\in\mathds{N}$ for each $a\in A$. A demanded balanced expansion ($A_{1},A_{2},y,f,D$) can be computed in $\mathcal{O}\left(|V|\,|E|\right)$ time. Furthermore, if there is an $A^{\prime}\subseteq A$ with $w(A^{\prime})+w(V(\mathcal{B}_{A^{\prime}}))\geq\sum_{a\in A^{\prime}}d_{a}$, then $A_{1}\neq\varnothing$.
The next graph structure that we use for our kernelization algorithms is introduced by Casel et al. [DBLP:conf/esa/Casel0INZ21] and is a combination of a balanced connected partition and a weighted crown decomposition, which is called balanced crown decomposition. Formally it is defined as follows (see also Figure 1 (right) for an illustration).
Definition 2.3 ($\lambda$-balanced crown decomposition).
A $\lambda$-balanced crown decomposition of a graph $G=(V,E,w)$ is a tuple $((C,H,\mathfrak{R},f))$, where $\{H,C,R\}$ is a partition of $V$, the set $\mathfrak{R}$ is a partition of $R$, and $f\colon\mathbb{CC}(G[C])\rightarrow H$, such that:
  • 1.
    N(C)H,
  • 2.
    w(Q)λ for each QCC(G[C]),
  • 3.
    f(Q)N(Q) for each QCC(G[C]),
  • 4.
    >+w(h)w(f-1(h))λ for each hH and
[…]
($\lambda,\lambda$)-CD
  • 5.
    $G[R^{\prime}]$ is connected and $\lambda<|R^{\prime}|\leq 3\lambda$ for each $R^{\prime}\in\mathfrak{R}$.
We use $\lambda$-BCD as an abbreviation for a $\lambda$-balanced crown decomposition. The authors [DBLP:conf/esa/Casel0INZ21] provide an algorithm that finds a $W$-BCD in polynomial time, as given in the following.
Theorem 2.4 (balanced crown decomposition theorem, [DBLP:conf/esa/Casel0INZ21] Theorem 2).
Let $G=(V,E,w)$ be a graph and $\lambda\in\mathds{N}$, such that each connected component in $G$ has weight larger than $\lambda$. A $\lambda$-balanced crown decomposition $(C,H,\mathfrak{R},f)$ of $G$ can be computed in $\mathcal{O}\left(r^{2}\,|V|\,|E|\right)$ time, where $r=|H|+|\mathfrak{R}|<|V|$ is at most the size of a maximum $(W+1)$-packing.
We end the preliminary section with a formal definition of the subgraph packing problem. Given a graph $G=(V,E)$ and two parameters $r,\lambda\in\mathds{N}$. We say that $P_{1},\dots,P_{m}\subseteq V$ is a $\lambda$-packing if for all $i,j\in[m]$ with $i\neq j$ the induced subgraph $G[P_{i}]$ is connected, $|P_{i}|\geq\lambda$, and $P_{i}\cap P_{j}=\varnothing$. The task is to find a $\lambda$-packing of size at least $r$.
3 Improved Kernels for VI, wVI and wCOC
The vertex integrity of a graph models finding an optimal balance between removing vertices and keeping small connected parts of a graph. Expressed more formally: Given a graph $G=(V,E)$ and a parameter $p\in\mathds{N}$. The task is to find a vertex set $S\subseteq V$ such that $|S|+\max_{Q\in\mathbb{CC}(G-S)}|Q|$ is at most $p$. In the weighted vertex integrity problem (wVI), i.e. the vertices have weights, the aim is that $w(S)+\max_{Q\in\mathbb{CC}(G-S)}w(Q)$ is at most $p$. We improve the vertex kernel of $p^{3}$ provided by Drange et al. [DBLP:journals/algorithmica/DrangeDH16] for both variants VI and wVI. To obtain such a kernel, they essentially established that vertices $v\in V$ with $w(v)+w(N(v))\geq p$ belong to a solution if we have a yes-instance in hand. A simple counting argument after removing those vertices provides a $p^{3}$ vertex kernel. We improve this outcome by employing crown decompositions, which offer valuable insights into the structural properties of vertex sets rather than focusing solely on individual vertices. Directly applying this method to problems like COC, as seen in prior works such as [DBLP:conf/esa/Casel0INZ21, DBLP:journals/tcs/ChenFSWY19, DBLP:conf/iwpec/KumarL16, DBLP:conf/tamc/XiaoK17], is challenging because we lack prior knowledge about the size of remaining components after solution removal. However, if we had at least a lower bound, we could theoretically safely reduce our instance accordingly. To establish such a bound, we engage in an interplay between packings and separators, where the balanced crown decomposition (BCD) proves instrumental. By identifying reducible structures in the input instance through embedding a DBE into BCD after determining a suitable bound, we prove the following theorems regarding VI and wVI.
Theorem 3.1.
The vertex integrity problem admits a vertex kernel of size $3p^{2}$ in time $\mathcal{O}\left(\log(p)|V|^{4}|E|\right)$.
Theorem 3.2.
The weighted vertex integrity problem admits a vertex kernel of size $3(p^{2}+p^{1.5}p_{\ell})$ in time $\mathcal{O}\left(\log(p)|V|^{4}|E|\right)$, where $p_{\ell}$ is at most the size of the largest component after removing a solution.
Closely related to wVI is the weighted component order connectivity problem (wCOC). Given a vertex-weighted graph $G=(V,E,w)$ and two parameters $k,W\in\mathds{N}$, the task is to find a vertex set $S\subseteq V$ such that $w(S)\leq k$ where each component weighs at most $W$. The techniques employed to derive the kernel for wVI can be seamlessly applied to wCOC, thereby enhancing the current state of the art vertex kernel of $kW(k+W)+k$. This kernel is also provided by Drange et al. [DBLP:journals/algorithmica/DrangeDH16] in a similar way as for wVI.
Theorem 3.3.
The weighted component order connectivity problem admits a vertex kernel of size $3\mu(k+\sqrt{\mu}W)$, where $\mu=\max(k,W)$. Furthermore, such a kernel can be computed in time $\mathcal{O}\left(r^{2}k|V|^{2}|E|\right)$, where $r$ is the size of a maximum $(W+1)$-packing.
Before we prove these theorems, we give some notations that we use in this section. We define them for the weighted case, as the unweighted case can be viewed in the same way with unit weights. We say that $S$ is a solution if it satisfies $w(S)+\max_{Q\in\mathbb{CC}(G-S)}w(Q)\leq p$ for wIV or $w(S)\leq k$ and $\max_{Q\in\mathbb{CC}(G-S)}w(Q)\leq W$ for wCOC. We denote an instance of wIV by $(G,p)$ and an instance of wCOC by $(G,k,W)$. We say that $(G,p)$ or $(G,k,W)$ is a yes-instance if there is a solution, otherwise, we say that it is a no-instance. Let $\mathcal{S}\subseteq 2^{V(G)}$ be all solutions in $(G,p)$. We define $p_{\ell}:=\min_{S\in\mathcal{S}}\left(\max_{Q\in\mathbb{CC}(G-S)}w(Q)\right)$, which is a lower bound for any solution $S\in\mathcal{S}$ on the size of the connected components after its removal. For all instances $(G,p)$ of wVI, we assume that $w(v)<p$ for every $v\in V$ and that $G$ contains a connected component of weight more than $p$. Otherwise, the instance is trivial.
3.1 Vertex Integrity
The rest of this section is dedicated to the proof of Theorem 3.1. First, we discuss the reducible structure we are looking for and then show how we can find it.
Lemma 3.4.
Let $(G,p)$ be an instance of VI and let $(C,H,f)$ be a $(p,p_{\ell})$-CD in $G$. Then, $(G,p)$ is a yes-instance if and only if $(G-\{H\cup C\},p-|H|)$ is a yes-instance.
In order to use Lemma 3.4, we must first determine a suitable lower bound for $p_{\ell}$. Additionally, we need to ensure the existence of a corresponding weighted crown decomposition when the input graph size exceeds $3p^{2}$, and that we can find it efficiently. This is where the BCD comes into play. For $\lambda\in\mathds{N}$ a $\lambda$-BCD $(C,H,\mathfrak{R},f)$ in $G$ can only be computed within the components of $G$ that have a size (or weight) greater than $\lambda$. For a better readability, we will consistently assume that when computing a $\lambda$-BCD, we disregard small components, meaning that if $(C,H,\mathfrak{R},f)$ is a $\lambda$-BCD for $G$, then $C\cup H\cup V(\mathfrak{R})$ are only the vertices that are contained in a component of size (or weight) more than $\lambda$ of $G$.
We specify the following result directly in the weighted version to also use it later.
Lemma 3.5.
Let $(G,p)$ be an instance of wVI and for $\lambda\in[p]$ let $(C,H,\mathfrak{R},f)$ be a $\lambda$-BCD in $G$. If $|H|+|\mathfrak{R}|>p$, then $\lambda<p_{\ell}$ for the instance $(G[C\cup H\cup V(\mathfrak{R})],p)$.
Note that the lower bound of $\lambda+1$ in Lemma 3.5 is applicable to the induced graph of $V^{\prime}=C\cup H\cup V(\mathfrak{R})$ within $G$. This implies that the isolated components of $G-V^{\prime}$, defined as having a size (or weight) of at most $\lambda$, can be safely removed. Their removal does not affect the decision problem of VI or wVI concerning $p$.
Clearly, a yes-instance cannot have a lower bound $p_{\ell}$ larger than $p$ as stated in the following corollary.
Corollary 3.6.
Let $(G,p)$ be an instance of wVI. If for a $p$-BCD $(C,H,\mathfrak{R},f)$ it holds that $|H|+|\mathfrak{R}|>p$, then $(G,p)$ is a no-instance.
The next lemma shows under which conditions we can find a $(p,\lambda)$-CD in $G$ with respect to a $\lambda$-BCD and the current graph size.
Lemma 3.7.
Let $(G,p)$ be an instance of the vertex integrity problem and for $\lambda\in[p]$ let $(C,H,\mathfrak{R},f)$ be a $\lambda$-BCD $(C,H,\mathfrak{R},f)$ in $G$ with $|H|+|\mathfrak{R}|\leq p$. If $|C|+|H|+|V(\mathfrak{R})|\geq 3p^{2}$, then $G^{\prime}=G[C\cup H]$ contains a $(p,\lambda)$-CD in $G$. Furthermore, we can extract it from $G^{\prime}$ in time $\mathcal{O}(|V(G^{\prime})||E(G^{\prime})|)\subseteq\mathcal{O}(|V(G)||E(G)|)$.
If we now combine Lemma 3.5 and Lemma 3.7, we can finally use Lemma 3.4 by finding a $\lambda\in[p-1]$ such that a ($\lambda+1$)-BCD $(C,H,\mathfrak{R},f)$ satisfies $|H|+|\mathfrak{R}|\leq p$, while a $\lambda$-BCD $(C^{\prime},H^{\prime},\mathfrak{R}^{\prime},f^{\prime})$ satisfies $|H^{\prime}|+|\mathfrak{R}^{\prime}|>p$, which only adds a factor of $\mathcal{O}(\log(p))$ to the computation cost. Formulated more precisely: For $Q\in\mathbb{CC}(C)$ we have $|Q|\leq\lambda+1\leq p_{\ell}$ as $\lambda<p_{\ell}$ by Lemma 3.5. Further, we remove all vertices that are in components of size at most $\lambda$, which is a save reduction rule as explained above. If the vertex size of the remaining input graph, i.e. $G[C\cup H\cup V(\mathfrak{R})]$ is at least $3p^{2}$, then we can find a ($p,p_{\ell}$)-CD for $G$ in $G[C\cup H]$ in polynomial time by Lemma 3.7, which is a reducible structure by Lemma 3.4.
With these facts in hand we can design a kernelization algorithm. It takes as input an instance $(G,p)$ of VI and returns an equivalent instance of with at most $3p^{2}$ vertices.
Find reducible structures (AlgVI):
  1. 1.
    Compute a $p$-BCD $(C_{1},H_{1},\mathfrak{R}_{1},f_{1})$. If $|H_{1}|+|\mathfrak{R}_{1}|>p$ return a trivial no-instance.
  2. 2.
    Compute a $1$-BCD $(C_{2},H_{2},\mathfrak{R}_{2},f_{2})$. Let $V^{\prime}=V(G)\setminus(C_{2}\cup H_{2}\cup V(\mathfrak{R}_{2}))$. If $|H_{2}|+|\mathfrak{R}_{2}|\leq p$ and $V^{\prime}\neq\varnothing$, then return $(G-V^{\prime},p)$. If $|H_{2}|+|\mathfrak{R}_{2}|\leq p$ and $V^{\prime}=\varnothing$, then compute a $(p,1)$-CD ($H_{2}^{\prime}$,$C_{2}^{\prime}$,$f_{2}^{\prime}$) in $G[H_{2}\cup C_{2}]$ and return $(G-\{H_{2}^{\prime}\cup C_{2}^{\prime}\},p-|H_{2}^{\prime}|)$.
  3. 3.
    Otherwise, for $\lambda\in[p-1]$ find a ($\lambda+1$)-BCD $(C,H,\mathfrak{R},f)$ and a $\lambda$-BCD $(C^{\prime},H^{\prime},\mathfrak{R}^{\prime},f^{\prime})$, such that $|H|+|\mathfrak{R}|\leq p$ and $|H^{\prime}|+|\mathfrak{R}^{\prime}|>p$. Let $V^{\prime\prime}=V(G)\setminus(C\cup H\cup V(\mathfrak{R}))$. If $V^{\prime\prime}\neq\varnothing$, then return $(G-V^{\prime\prime},p)$. Otherwise, compute a $(p,\lambda+1)$-CD ($H^{\prime\prime}$,$C^{\prime\prime}$,$f^{\prime\prime}$) in $G[C\cup H]$ for $G$ and return $(G-\{C^{\prime\prime}\cup H^{\prime\prime}\},p-|H^{\prime\prime}|)$.
We already explained why we can safely remove $V^{\prime}$ or $V^{\prime\prime}$ in step 2 and 3 if these are not empty. Note that if $V^{\prime}$ or $V^{\prime\prime}$ is empty, then the vertex sets of the corresponding BCD’s, i.e. $C_{2},H_{2},V(\mathfrak{R}_{2})$ and $C,H,V(\mathfrak{R})$, respectively, form a partition of $V(G)$. The correctness of step 1 is implied by Corollary 3.6. Step 2 is correct because $p_{\ell}\geq 1$ and if $V^{\prime}=\varnothing$ we obtain by Lemma 3.7 and $|V|=|C_{2}|+|H_{2}|+|V(\mathfrak{R}_{2})|\geq 3p^{2}$ that there must be a $(p,1)$-CD if $|H_{2}|+|\mathfrak{R}_{2}|\leq p$. Analogously, for step 3 as $p_{\ell}\geq\lambda+1$ by Lemma 3.5 in this step. For the overall correctness of algorithm AlgVI, it remains to be proven that if we reach step 3, the required $\lambda$ exists. For a $\lambda$-BCD in a binary search, we cannot ensure that $|H|+|\mathfrak{R}|$ increases with decreasing $\lambda$ values, because the sizes of $H$ and $\mathfrak{R}$ are not necessarily monotonic with respect to $\lambda$. Note, for example, that the elements in $\mathfrak{R}$ have a size range from $\lambda+1$ to $3\lambda$. For a successful binary search, however, it is sufficient to know that there is a $\lambda$ that satisfies the desired properties for step 3. Since we only enter this step if the extreme cases ($p$-BCD and 1-BCD) in step 1 and 2 hold, there has to exist at least one such $\lambda$ value in-between.
Finally, we prove the specified running time of Theorem 3.1, which completes the proof of Theorem 3.1. Note that for a yes-instance we have to call the algorithm AlgVI at most $|V|$ times to guarantee the desired kernel.
Lemma 3.8.
Algorithm AlgVI runs in time $\mathcal{O}(\log(p)|V|^{3}|E|)$.
3.2 Weighted Vertex Integrity and Component Order Connectivity
In this section, we shift our focus to the weighted variants, namely wVI and wCOC. While utilizing the packing size of the associated BCD offers a starting point for deriving a lower bound for VI, it proves insufficient for improvements in the weighted setting. Therefore, we integrate the weight of the separator $H$ within a BCD $(C,H,\mathfrak{R},f)$ into our analysis. Additionally, we incorporate two distinct DBE’s into a BCD in case a reduction is not achievable within the respective setting. This approach enables us to establish a tighter lower bound for $p_{\ell}$, estimate the remaining instance size more accurately to obtain the desired kernelization results (cf. Theorems 3.2 and 3.3), or a identify no-instance.
For improved readability, we extend the notation of the inverse function $f^{-1}$. Henceforth, sometimes it returns a vertex set that is the union of the corresponding vertex sets, which is always clear from the context. For instance, for $h\in H$ or $H^{\prime}\subseteq H$ we have $f^{-1}(h)=V(f^{-1}(h))$ and $f^{-1}(H^{\prime})=V(f^{-1}(H^{\prime}))$, respectively. Let us start by explaining the type of reducible structure we are searching for.
Lemma 3.9.
Let $G=(V,E,w)$ be a vertex weighted graph, $a,b\in\mathds{N}$, $(H,H^{\prime},b,f,D)$ a DBE in $G$ with $D=\{d_{h}\}_{h\in H}$, where $d_{h}=a-2+b\cdot(w(h)+1)$ for each $h\in H$, and let $C=f^{-1}(H)$.
  1. 1.
    Let $(G,p)$ be an instance of the weighted vertex integrity problem, $a=p$ and $1\leq b\leq p_{\ell}$. Then, $(G,p)$ is a yes-instance if and only if $(G-\{H\cup C\},p-w(H))$ is a yes-instance.
  2. 2.
    Let $(G,k,W)$ be an instance of the weighted component order connectivity problem and $a,b=W$. Then, $(G,k,W)$ is a yes-instance if and only if $(G-\{H\cup C\},k-w(H),W)$ is a yes-instance.
We are now introducing the two DBE’s mentioned in conjunction with a BCD. Let $G=(V,E,w)$ be a vertex weighted graph, $a,s,\lambda\in\mathds{N}$, $(C,H,\mathfrak{R},f)$ a $\lambda$-BCD with $\lambda\in[a]$ and $|H|+|\mathfrak{R}|\leq s$, where $\max_{v\in V}w(v)\leq\max(a,s)=:\mu$. Let $D^{Y}=\{d^{Y}_{h}\}_{h\in H}$ and $D^{Z}=\{d^{Z}_{h}\}_{h\in H}$ be demands, where $d^{Y}_{h}=a-2+\lambda\cdot(w(h)+1)$ and $d^{Z}_{h}=w(h)-1+(\sqrt{s}+1)\lambda$ for each $h\in H$. Let $(Y_{1},Y_{2},\lambda,f_{Y},D^{Y})$ and $(Z_{1},Z_{2},\lambda,f_{Z},D_{Z})$ be DBE’s in $G[C\cup H]$ with $Y_{1},Y_{2},Z_{1},Z_{2}\subseteq H$ and $f_{Y}^{-1}(Y_{1}),f_{Y}^{-1}(Y_{1}),f_{Z}^{-1}(Z_{1}),f_{Z}^{-1}(Z_{2})\subseteq\mathbb{CC}(C)$. Observe that if $Y_{1}\neq\varnothing$ (resp. $Z_{1}\neq\varnothing$) then this set separates $f^{-1}_{Y}(Y_{1})$ (res. $f^{-1}_{Z}(Z_{1})$) from the rest of the graph, since $H$ separates $C$ from the rest of the graph. Thus, if $a=p$ and $\lambda\leq p_{\ell}$ considering wVI or $a,\lambda=W$ considering wCOC and $Y_{1}\neq\varnothing$, then in both cases we would have a reducible structure for the corresponding problem (cf. Lemma 3.9). We conclude this section by presenting two crucial lemmas (the continuation can be found in LABEL:appendix::continuation-wVI-wCOC). These lemmas serve as the foundation for designing the algorithms needed to prove Theorems 3.2 and 3.3. The first lemma aims to estimate the instance size, while the second lemma helps establish a more precise lower bound for $p_{\ell}$ or identify instances with no feasible solutions.
Lemma 3.10.
Let $w(Z_{1})\leq 2\mu^{1.5}$. If $Y_{1}=\varnothing$, then $w(V)<3\mu(s+\sqrt{\mu}\lambda)$.
Lemma 3.11.
If $w(Z_{1})>2\mu^{1.5}$, then there is no separator $S$ such that $w(S)\leq s$ and $\max_{Q\in\mathbb{CC}(G-S)}w(Q)\leq\lambda$.
4 Kernels for Component Order Connectivity
In this section, we consider COC, a refined version of VI. Given a graph $G=(V,E)$ and parameters $k,W\in\mathds{N}$, COC aims to remove at most $k$ vertices so that resulting components have at most $W$ vertices each. We present a kernelization algorithm that provides a $2kW$ vertex-kernel in an FPT-runtime when parameterized by the size of a maximum ($W+1$)-packing, as stated in the following theorem.
Theorem 4.1.
A vertex kernel of size $2kW$ for the component order connectivity problem can be computed in time $\mathcal{O}(r^{3}|V||E|\cdot r^{\min(3r,k)})$, where $r\leq k$ is the size of a maximum $(W+1)$-packing.
A kernel of size $2kW$ is also presented in [DBLP:conf/iwpec/KumarL16] with an exponential runtime using linear programming methods, which however, as mentioned above, can be improved to an FPT-runtime in the parameter $W$. In contrast, our result is fully combinatorial and has an FPT-runtime in the parameter of a maximum $(W+1)$-packing $r\leq k$. Moreover, we showcase how our algorithm transforms into a polynomial one for two cases: when $W=1$ (Vertex Cover) and for claw-free graphs. While generalizing our FPT-runtime algorithm into polynomial time eludes us, our understanding of crown decompositions structural properties holds promise for future progress.
Before, i.e. for VI, wVI, and wCOC, crown structures were integrated into the head and crown of the BCD. In contrast, we now incorporate crown structures into the body of the BCD. Of particular interest are the so-called strictly reducible pairs introduced by Kumar and Lokshtanov [DBLP:conf/iwpec/KumarL16] who prove that such a structure must exist in graphs with more than $2kW$ vertices.
Definition 4.2 ((strictly) reducible pair).
For a graph $G=(V,E)$ and a parameter $W\in\mathds{N}$, a pair $(A,B)$ of vertex disjoint subsets of $V$ is a reducible pair for COC if the following conditions are satisfied:
  • $N(B)\subseteq A$.
  • The size of each $Q\in\mathbb{CC}(G[B])$ is at most $W$.
  • There is an assignment function $g\colon\mathbb{CC}(G[B])\times A\to\mathds{N}_{0}$, such that
    • for all $Q\in\mathbb{CC}(G[B])$ and $a\in A$, if $g(Q,a)\neq 0$, then $a\in N(Q)$
    • for all $a\in A$ we have $\sum_{Q\in\mathbb{CC}(G[B])}g(Q,a)\geq 2W-1$,
    • for all $Q\in\mathbb{CC}(G[B])$ we have $\sum_{a\in A}g(Q,a)\leq|Q|$.
In addition, if there exists an $a\in A$ such that $\sum_{Q\in\mathbb{CC}(G[B])}g(Q,a)\geq 2W$, then $(A,B)$ is a strictly reducible pair.
We say $(A,B)$ is a minimal (strictly) reducible pair if there does not exist a (strictly) reducible pair $(A^{\prime},B^{\prime})$ with $A^{\prime}\subset A$ and $B^{\prime}\subseteq B$. A (strictly) reducible pair is basically a weighted crown decomposition where $A$ is the head and $B$ is the crown.
In essence, the kernelization algorithm outlined below focuses on analyzing pairs $(A,B)$ in a $W$-BCD. It turns out that the head vertices $A$ have specific traits in a $W$-BCD, making it possible to locate them. Structurally, the algorithm operates as a bounded search tree, hence it is presented recursively. It takes an instance $(G,k,W)$ of COC as input, where $|V(G)|>2kW$ and each component of $G$ has at least $W+1$ vertices. The size limits for the input graph are not arbitrary; if $|V(G)|\leq 2kW$, there is nothing to do, and removing components smaller than $W+1$ is a safe reduction.
For $V^{\prime}\subseteq V$ we define $\texttt{sep}_{W}(V^{\prime})\in\mathds{N}$ as the cardinality of a minimum $W$-separator in $G[V^{\prime}]$, and $\texttt{arg\hskip 2.84544ptsep}_{W}(V^{\prime})\subset V$ as an argument suitable to this cardinality. For a graph $G^{\prime}\subseteq G$ let $G^{>W}(G^{\prime})$ be the graph obtained by removing all components of size at most $W$ from $G$. Lastly, for a $W$-BCD $(C,H,\mathfrak{R},f)$ we define :=$\mathfrak{R}^{\prime}:=\{R\in\mathfrak{R}\mid|R|>2W\text{ and }\texttt{sep}_{W}(R)=1\}$ and $S_{\mathfrak{R}^{\prime}}:=\bigcup_{R\in\mathfrak{R}^{\prime}}\texttt{arg\hskip 2.84544ptsep}_{W}(R)$. (We show in LABEL:lemma::UniqueSep the uniqueness of $\texttt{arg\hskip 2.84544ptsep}_{W}(R)$ for $R\in\mathfrak{R}^{\prime}$).
Find reducible structures (
  1. 1.
    Compute a a $W$-BCD $(C,H,\mathfrak{R},f)$ in $G$ and initialize $t=|H|+|\mathfrak{R}|$ and $S=\varnothing$.
  2. 2.
    Let $G^{\prime}=G^{>W}(G-S)$. If $G^{\prime}$ is an empty graph return a trivial yes-instance. Otherwise, compute a $W$-BCD $(C,H,\mathfrak{R},f)$ in $G^{\prime}$.
    1. (a)
      If $|H|+|\mathfrak{R}|>k$ return a trivial no-instance.
    2. (b)
      Let $\mathcal{Q}$ be the connected components of size at most $W$ in $G-S$. Let $A=S\cup H$ and $B=C\cup V(\mathcal{Q})$. Compute a DBE $(A_{1},A_{2},W,f,2W-1)$ in $G[A\cup B]$ by using Theorem 2.2. If $A_{1}\neq\varnothing$, then terminate the algorithm and return $(G-(A_{1}\cup f^{-1}(A_{1})),k-|A_{1}|,W)$.
  3. 3.
    If the depth of the recursion is more than $\min(3t,k)$, then return.
  4. 4.
    For each $v\in H\cup S_{\mathfrak{R}^{\prime}}$:
    • Add $v$ to $S$ and recurse from step 2.
  5. 5.
    Return a trivial no-instance.
The interesting part of the algorithm is the localization of a minimal strictly reducible pair if it exists in the graph. Let $(A,B)$ be a minimal strictly reducible pair in $G$. As already mentioned, algorithm
Algorithm 2.
2kw is basically a bounded search tree with a working vertex set $S$, which are potential head vertices. The crucial step is ensuring that $S$ consistently matches the vertex set $A$. Once this alignment is achieved, localization of $(A,B)$ is assured, as step 2b (specifically Theorem 2.2) guarantees its extraction.
We conclude this part of the section by presenting a crucial lemma along with its implication for a depth-wise progression within the bounded search tree towards $S=A$ in algorithm