### v2

Abstract
Over the past decades, various metrics have emerged in graph theory to grasp the complex nature of network vulnerability. In this paper, we study two specific measures: (weighted) vertex integrity (wVI) and (weighted) component order connectivity (wCOC). These measures not only evaluate the number of vertices that need to be removed to decompose a graph into fragments, but also take into account the size of the largest remaining component. The main focus of our paper is on kernelization algorithms tailored to both measures. We capitalize on the structural attributes inherent in different crown decompositions, strategically combining them to introduce novel kernelization algorithms that advance the current state of the field. In particular, we extend the scope of the balanced crown decomposition provided by Casel et al. [DBLP:conf/esa/Casel0INZ21] and expand the applicability of crown decomposition techniques.
In summary, we improve the vertex kernel of VI from $p^{3}$ to $3p^{2}$, and of wVI from $p^{3}$ to $3(p^{2}+p^{1.5}p_{\ell})$, where $p_{\ell}<p$ represents the weight of the heaviest component after removing a solution. For wCOC we improve the vertex kernel from $\mathcal{O}(k^{2}W+kW^{2})$ to $3\mu(k+\sqrt{\mu}W)$, where $\mu=\max(k,W)$. We also give a combinatorial algorithm that provides a $2kW$ vertex kernel in fixed-parameter tractable time when parameterized by $r$, where $r\leq k$ is the size of a maximum $(W+1)$-packing. We further show that the algorithm computing the $2kW$ vertex kernel for COC can be transformed into a polynomial algorithm for two special cases, namely when $W=1$, which corresponds to the well-known vertex cover problem, and for claw-free graphs. In particular, we show a new way to obtain a $2k$ vertex kernel (or to obtain a 2-approximation) for the vertex cover problem by only using crown structures.
keywords:
Crown Decomposition, Kernelization, Vertex Integrity, Component Order Connectivity
1 Introduction
In the study of graph theory different scales have emerged over the past decades to capture the complex nature of network vulnerability. While the main focus is on the connectivity of vertices and edges, subtle aspects of vulnerability such as the number of resulting components, the size distribution of the remaining components, and the disparity between them are becoming increasingly interesting [DBLP:conf/esa/BentertHK23, DBLP:journals/tcs/GimaHKKO22, DBLP:conf/walcom/GimaHKM0O24, DBLP:journals/corr/abs-2402-09971, DBLP:journals/algorithmica/GimaO24, DBLP:conf/isaac/LampisM21]. Our focus in this study is on two specific ways to measure vulnerability: (weighted) vertex integrity and (weighted) component order connectivity. These measures not only evaluate the number of vertices that need to be removed to decompose a graph into fragments, but also take into account the size of the largest remaining component. Incorporating these aspects provides a more comprehensive understanding of network resilience.
Informally, vertex integrity (VI) is a model for the balance between removing few vertices and keeping small connected parts of a graph. More formally, given a graph $G=(V,E)$ and a number $p\in\mathds{N}$, the task for VI is to find a set of vertices $S\subseteq V$ such that $|S|$ plus the size of the largest component when removing $S$ from $G$ is at most $p$. In the vertex weighted version (wVI), the goal is bounding the total weight of the removed vertices plus the weight of the heaviest component by $p$. This problem was introduced by Barefoot et al. [barefoot1987vulnerability] as a way to measure vulnerability of communication networks. Recently, it has drawn the interest of the parameterized complexity community due to its status as a natural parameter that renders numerous NP-hard problems amenable to fixed parameter tractability (FPT). This means that for these problems, solutions can be computed within a time frame represented by $f(p)\cdot n^{\mathcal{O}(1)}$, where $f$ is a computable function [DBLP:conf/isaac/LampisM21]. It is interesting to see how vertex integrity relates to other well-known measures of network structure. It imposes greater constraints compared to metrics such as treedepth, treewidth, or pathwidth, as the vertex integrity of a graph serves as an upper bound for these parameters. However, it encompasses a wider range of scenarios compared to vertex cover, where a vertex cover of a graph is an upper bound for its vertex integrity. This makes it a key in understanding how to efficiently solve problems in the world of network analysis.
The measure component order connectivity (COC) can be seen as the refined version of VI. Given a graph $G=(V,E)$ and two parameters $k,W\in\mathds{N}$, the goal of COC is to remove $k$ vertices such that each connected component in the resulting graph has at most $W$ vertices — also known in the literature as the $W$-separator problem or $\alpha$-balanced separator problem, where $\alpha\in(0,1)$ and $W=\alpha|V|$. In the vertex-weighted version (wCOC), the goal is to remove vertices of total weight at most $k$ such that the weight of the heaviest remaining component is at most $W$. An equivalent view of this problem is to search for the minimum number of vertices required to cover or hit every connected subgraph of size $W+1$. In particular, $W=1$ corresponds to covering all edges, showing that the COC is a natural generalization of the vertex cover problem.
The focus of the paper is on kernelization algorithms tailored for both weighted and unweighted versions of VI and COC when parameterized by $p$ and $k+W$, respectively. Kernelization algorithms can be thought of as formalized preprocessing techniques aimed to reduce insances of optimization problems. Of particular interest in this work are crown decompositions, which are generally used as established structures for safe instance reduction — where “safe” means that any optimal solution to the reduced instance can efficiently be transformed into an optimal solution of the original one. Essentially, a crown decomposition partitions the vertex set into three disjoint components: the crown, the head, and the body. Here, the head acts as a separator between the crown and the body. This structural arrangement becomes useful when specific relationships between the head and the crown are required. Such relationships ultimately enable us to shrink instances by eliminating these designated parts from the graph. The properties of this structural layout, coupled with its existence depending on the instance size, enable the development of efficient kernelization algorithms for different problem domains. Notably, crown decompositions have also recently found utility in approximation algorithms for graph packing and partitioning problems [DBLP:conf/esa/Casel0INZ21]. For further exploration of crown decompositions, including their variations and applications, we recommend the comprehensive survey paper by Jacob et al. [DBLP:journals/algorithms/JacobMR23].
Our methods take advantage of the structural characteristics found in various crown decompositions, leading to the development of new kernelization algorithms that improve the state of the art. In essence, this work expands upon the applications of the balanced crown decomposition introduced by Casel et al. [DBLP:conf/esa/Casel0INZ21], specifically by integrating different crown decompositions into this framework.
Related Work
Recently, vertex integrity (VI) has received considerable attention as a structural graph parameter [DBLP:conf/esa/BentertHK23, DBLP:journals/tcs/GimaHKKO22, DBLP:journals/algorithmica/GimaO24, DBLP:conf/isaac/LampisM21]. Some results in the literature are stated in terms of the fracture number of the graph [DBLP:journals/algorithmica/GanianOR21], which is a closely related parameter that takes the maximum of the modulator size and the largest remaining component; in particular, fracture number is always within a factor of two from vertex integrity. Gima et al. [DBLP:conf/walcom/GimaHKM0O24] conducted a systematic investigation into structural parameterizations of computing the VI and wVI which was further extended by Hanaka et al. [DBLP:journals/corr/abs-2402-09971]. Additionally, there are notable results concerning special graph classes [DBLP:journals/dam/BaggaBGLP92, clark1987computational, DBLP:journals/algorithmica/DrangeDH16, DBLP:journals/dam/KratschKM97, DBLP:journals/ijcm/LiZZ08]. In our context, regarding related work on VI and wVI, Fellows and Stueckle presented an algorithm that solves the problem in time $\mathcal{O}(p^{3p}n)$ [fellows1989immersion]. This was subsequently improved by Drange et al. [DBLP:journals/algorithmica/DrangeDH16], even for the weighted case, to $\mathcal{O}(p^{p+1}n)$. In the same paper, they presented the first vertex-kernel of size $p^{3}$ for both VI and wVI.
Considering, COC and wCOC, it is unlikely that kernelization algorithms for these problems can be achieved by considering $k$ or $W$ alone in polynomial time. Indeed, $W=1$ corresponds to the NP-hard vertex cover problem, which shows that $W$ (alone) is not a suitable parameter. For the parameter $k$, the problem is $W[1]$-hard even when restricted to split graphs [DBLP:journals/algorithmica/DrangeDH16]. These lower bounds lead to the study of parameterization by $k+W$. The best known algorithm with respect to these parameters finds a solution in time $n^{O(1)}\cdot 2^{\mathcal{O}(\log(W)\cdot k)}$ [DBLP:journals/algorithmica/DrangeDH16]. Unless the exponential time hypothesis fails, the authors prove that this running time is tight in the sense that there is no algorithm that solves the problem in time $n^{\mathcal{O}(1)}\cdot 2^{o(\log(W)\cdot k)}$. The best known approximation algorithm has a multiplicative gap guarantee of $\mathcal{O}(\log(W))$ to the optimal solution with a running time of $n^{\mathcal{O}(1)}\cdot 2^{\mathcal{O}(W)}$ [DBLP:journals/corr/Lee16c]. In [DBLP:journals/corr/Lee16c], the authors also showed that the superpolynomial dependence on $W$ may be needed to achieve a polylogarithmic approximation. Using this algorithm as a subroutine, the vertex integrity can be approximated within a factor of $\mathcal{O}(\log(\text{OPT}))$, where OPT is the vertex integrity.
Regarding kernelization algorithms, there is a sequence of results which successively improve the vertex-kernel of COC. The first results came from Chen et al. [DBLP:journals/tcs/ChenFSWY19] and Drange et al. [DBLP:journals/algorithmica/DrangeDH16], who provided kernels of size in $\mathcal{O}(kW^{3})$ and $\mathcal{O}(k^{2}W+kW^{2})$, respectively. The result of Drange et al. also holds for wCOC and is the only result for this case. This was improved simultaneously by Xiao [DBLP:journals/jcss/Xiao17a] as well as by Kumar and Lokshtanov [DBLP:conf/iwpec/KumarL16] to a $\mathcal{O}(kW^{2})$ kernel. These works also provide the first $\mathcal{O}(kW)$ kernels, but with different constants and running times. Kumar and Lokshtanov [DBLP:conf/iwpec/KumarL16] present a $2kW$ kernel in a running time of $n^{\mathcal{O}(W)}$ by using linear programming (LP) methods with an exponential number of constraints. The runtime can be improved to $2^{\mathcal{O}(W)}\cdot n^{\mathcal{O}(1)}$ as already mentioned in the book of Fomin et al. [fomin2019kernelization] (Section 6.4.2). Roughly speaking, the idea is to use the ellipsoid method with separation oracles to solve the linear program, where the separation oracle uses a method called color coding to find violated constraints that makes it polynomial in $W$. Note that if $W$ is a constant this $2kW$ kernel is a polynomial time kernel improving on some previous results. This includes for instance the improvement of the $5k$ kernel provided by Xiao and Kou [DBLP:conf/tamc/XiaoK17] to a $4k$ kernel for the well-studied $P_{2}$-covering problem, where a $P_{2}$ is a path with 2 edges. The first linear kernel in both parameters in polynomial time, i.e. an $\mathcal{O}(kW)$ vertex kernel, is presented by Xiao [DBLP:journals/jcss/Xiao17a], who provides a $9kW$ vertex kernel. Finally, this was improved by Casel et al. [DBLP:conf/esa/Casel0INZ21] to a $3kW$ vertex kernel, which also holds for a more general setting. Namely, to find $k$ vertices in a vertex weighted graph such that after their removal each component weighs at most $W$. Note that the weights of the separator, i.e. the chosen $k$ vertices, play no role in this problem compared to wCOC. With the exception of the $\mathcal{O}(k^{2}W+kW^{2})$ vertex kernel of Drange et al. [DBLP:journals/algorithmica/DrangeDH16], all achieved vertex kernels essentially use crown structures.
Our Contribution
The provided running times are based on an input graph $G=(V,E)$. We improve the vertex kernel for VI from $p^{3}$ to $3p^{2}$ in time $\mathcal{O}\left(\log(p)|V|^{4}|E|\right)$. For wVI, we improve it to $3(p^{2}+p^{1.5}p_{\ell})$ in time $\mathcal{O}\left(\log(p)|V|^{4}|E|\right)$, where $p_{\ell}\leq p$ represents the weight of the heaviest component after removing a solution.
To better explain the results of COC and wCOC, consider the problem of a maximum $\lambda$-packing for $\lambda\in\mathds{N}$. Given a (vertex weighted) graph $G$, this problem aims to maximize the number of disjoint connected subgraphs, each of size (or weight) at least $\lambda$. It is worth noting that the size of a maximum $(W+1)$-packing serves as a lower bound on the size of an optimal solution of COC (or wCOC), since each element in the packing must contain at least one vertex of it — in terms of linear programming, it is the dual of COC.
For wCOC we improve the vertex kernel of $\mathcal{O}(k^{2}W+kW^{2})$ to $3\mu(k+\sqrt{\mu}W)$ in time $\mathcal{O}\left(r^{2}k|V||E|\right)$, where $\mu=\max(k,W)$ and $r\leq|V|$ is the size of a maximum $(W+1)$-packing. For the unweighted version, we provide a $2kW$ vertex kernel in an FPT-runtime of $\mathcal{O}(r^{3}|V||E|\cdot r^{\min(3r,k)})$, where $r\leq k$ is the size of a maximum $(W+1)$-packing. Comparing this result with the state of the art, disregarding the FPT-runtime aspect, we improve upon the best-known polynomial algorithm, achieving a kernel of size $3kW$ [DBLP:conf/esa/Casel0INZ21]. A $2kW$ vertex kernel is also presented in [DBLP:conf/iwpec/KumarL16], albeit with an exponential runtime using linear programming methods, which, as mentioned, can be enhanced to an FPT-runtime regarding parameter $W$. In contrast, our result is entirely combinatorial and has an FPT-runtime in the parameter of a maximum $(W+1)$-packing $r\leq k$. It should be noted that, strictly speaking, the $2kW$ vertex kernel of [DBLP:conf/iwpec/KumarL16] and our work cannot be considered a kernel, given that the runtime dependency is exponential with respect to the parameters $W$ and $k$, respectively. However, for the sake of simplicity, we refer to it as a kernel, with an explicit mention of the runtime dependency. As previously stated, note that it is unlikely that an FPT-runtime will be sufficient for solving COC (or wCOC) when considering either $W$ or $k$ alone. We further show that the algorithm computing the $2kW$ vertex kernel for COC can be transformed into a polynomial algorithm for two special cases. The first case arises when $W=1$, i.e., for the vertex cover problem. Here, we provide a new method for obtaining a vertex kernel of $2k$ (or obtaining a $2$-approximation) using only crown decompositions. The second special case is for the restriction of COC to claw-free graphs. Notably, it remains open whether there is a polynomial-time algorithm for COC on claw-free graphs, or whether the problem is NP-hard on this class; we defer this question to future work.
Regarding these special cases, until 2017, a $2k$ vertex kernel for $W=1$ was known through crown decompositions, albeit computed using both crown decompositions and linear programming. Previously, only a $3k$ vertex kernel was known using crown decompositions alone. In 2018, Li and Zhu [DBLP:journals/tcs/LiZ18] provided a $2k$ vertex kernel solely based on crown structures. They refined the classical crown decomposition, which possessed an additional property allowing the remaining vertices of the graph to be decomposed into a matching and odd cycles after exhaustively applying the corresponding reduction rule. In contrast, our algorithm identifies the reducible structures, which must exist if the size of the input graph exceeds $2k$. For claw-free graphs, we are not aware of previous work regarding kernels for VI or COC on this class. However, the classical result of Minty [DBLP:journals/jct/Minty80] states that one can find vertex cover of smallest weight in a claw-free graph in polynomial time.
Unfortunately, we were unable to transform the FPT-runtime kernelization algorithm into a polynomial-time algorithm in general. Nevertheless, we believe that our insights into the structural properties of crown decompositions pave the way to achieving this goal.
2 Preliminaries
In the upcoming paragraphs, we discuss the standard terminology related to graphs and parameterized complexity. Additionally, we introduce some crown structures that are utilized throughout this paper.
Graph Terminology
Let $G=(V,E)$ be a graph. We use $V(G)$ and $E(G)$ to denote $V$ and $E$, respectively. We define the size of a subgraph $G^{\prime}\subseteq G$ as the number of its vertices, where we denote the size of $G$ by $n$. For $v\in V(G)$ we denote its neighborhood by $N(v):=\{u\in V(G):uv\in E(G)\}$. Let $V^{\prime}\subseteq V$ be a vertex subset. We define $G[V^{\prime}]$ as the induced subgraph of $V^{\prime}$, $G-V^{\prime}:=G[V\setminus V^{\prime}]$ and $N(V^{\prime}):=\left(\bigcup_{v\in V^{\prime}}N(v)\right)\setminus V^{\prime}$. For a vertex $v$ we define $G+v$ as the graph $G$ with an additional isolated vertex $v$. Similarly, for a graph $H$ we define $G+H$ as the graph $G$ with an additional isolated graph $H$. For sets of vertex sets $\mathcal{V}\subseteq 2^{V}$ we abuse notation and use $V(\mathcal{V})=\bigcup_{Q\in\mathcal{V}}Q$ and $N(\mathcal{V}):=N(V(\mathcal{V}))=\left(\bigcup_{Q\in\mathcal{V}}N(Q)\right)\setminus V(\mathcal{V})$. We define $\mathbb{CC}(G)\subset 2^{V}$ as the connected components of $G$ as vertex sets. For vertex subsets $H,C\subset V$, we define a function $f\colon\mathbb{CC}(C)\to H$. For any subset $H^{\prime}\subseteq H$, we frequently use the inverse $f^{-1}(H^{\prime})$, which returns sets of vertices. Typically, however, we are interested in the union of these vertex sets, denoted $V(f^{-1}(H^{\prime}))$. For readability, we may omit the vertex operator when it is clear from the context that we refer to the union. We use the same terminology for vertex weighted graphs, denoted by $G=(V,E,w)$ where $w$ is a weight function $w\colon V\to\mathds{N}$. Let $G=(V,E,w)$ be a vertex weighted graph. For $V^{\prime}\subseteq V$, $G^{\prime}\subseteq G$ (meaning that $G^{\prime}$ is a subgraph of $G$) and $\mathcal{V}\subset 2^{V}$ we define $w(V^{\prime})$, $w(G^{\prime})$ and $w(G^{\prime})$ as $\sum_{v\in V^{\prime}}w(v)$, $\sum_{v\in V(G^{\prime})}w(v)$ and $\sum_{v\in V(\mathcal{V})}w(v)$, respectively.
Parameterized Terminology
We use the standard terminology for parameterized complexity, which is also used, for example, in [downey2012parameterized, fomin2019kernelization]. A parameterized problem is a decision problem with respect to certain instance parameters. Instances of a parameterized problem are usually given as tuple $(I,k)$, where $k$ is the parameter.
If there exists an algorithm that decides each instance $(I,k)$ of a parameterized problem $\Pi$ in time $f(k)\cdot|I|^{c}$, where $f$ is a computable function and $c$ is a constant, then $\Pi$ is called fixed-parameter tractable. We say $(I,k)$ is a yes-instance if the answer to the decision problem is positive, otherwise we say $(I,k)$ is a no-instance.
Of particular interest in this work are kernelizations, which can be roughly described as formalized preprocessing. More formally, a polynomial algorithm is called a kernelization for a parameterized problem $\Pi$ if it maps any instance $(I,k)$ of $\Pi$ to an instance $(I^{\prime},k^{\prime})$ of $\Pi$ such that $(I^{\prime},k^{\prime})$ is a yes-instance if and only if $(I,k)$ is a yes-instance, $|I^{\prime}|\leq g(k)$, and $k^{\prime}\leq g^{\prime}(k)$ for computable functions $g,g^{\prime}$.
Crown Decompositions
We present a more general variant of crown decomposition that also captures commonly used crown decompositions or expansions (strictly speaking the relevant parts of it), which we explain in a moment. This novel decomposition can be easily derived from existing results of [DBLP:conf/esa/Casel0INZ21], but to the best of our knowledge has never been used in this form before (see Fig. 1 (left) for an illustration).
Definition 1 (Demanded balanced expansion and weighted crown decomposition).
For a graph $G=(A\cup B,E,w)$, a partition $A_{1}\cup A_{2}$ of $A$, a function $f\colon\mathbb{CC}(G[B])\to A$, demands $D=\{d_{a}\}_{a\in A}$ with $d_{a}\in\mathds{N}$ for each $a\in A$ and $y\in\mathds{N}$ the tuple $(A_{1},A_{2},y,f,D)$ is a demanded balanced expansion (DBE) if
  1. 1.
    $w(Q)\leq y$ for each $Q\in\mathbb{CC}(G[B])$
  2. 2.
    $f(Q)\in N(Q)$ for each $Q\in\mathbb{CC}(G[B])$
  3. 3.
    $N(f^{-1}(A_{1}))\subseteq A_{1}$
  4. 4.
    $w(a)+w\left(f^{-1}(a)\right)\begin{cases}>d_{a}-y+1\text{ for each }a\in A_{1}\\ \leq d_{a}+y-1\text{ for each }a\in A_{2}\end{cases}$
To simplify the notation we introduce two further special cases of a DBE:
  • If the demands are the same for each $a\in A$, e.g. $d_{a}=x$ for every $a\in A$ with $x\in\mathds{N}$, then we write only the value $x$ instead of a vector $D=\{d_{a}\}_{a\in A}$ in a DBE-tuple, i.e. $(A_{1},A_{2},y,f,D)=(A_{1},A_{2},y,f,x)$.
  • For $q\in\mathds{N}$ we call $(C,H,f)$ a ($q,y$) crown decomposition (($q,y$)-CD) if it is a $(H,\varnothing,y,f,q+y-1)$ DBE with $H=A$ and $C=B=f^{-1}(H)$. (This term simplifies the reference to the reducible structure of crown $C$ and head $H$.)
[…]
Figure 1:
Left: Let $A=\{a_{1},a_{2},a_{3},a_{4}\}$ be ordered in a top down manner, $w(Q)=1$ for every $Q\in\mathbb{CC}(G[B])$ and $w(a)=1$ for every $a\in A$. Then, ($(\{a_{1},a_{2}\},\{a_{3},a_{4}\},1,f,\{3,1,3,4\})$ is a DBE, where the assignment $f$ are depicted with corresponding colored bold edges.
Right: A $\lambda$-balanced crown decomposition, where the assignment $f$ are depicted with corresponding colored bold edges. The two dashed lines illustrate that $w(h)+w(f^{-1}(h))>\lambda$ for every $h\in H$ while $w(Q)\leq\lambda$ for every $Q\in\mathbb{CC}(G[C])$.
Considering a graph instance $G=(V,E)$, e.g. from VI or COC, Definition 1 usually describes only a subgraph where the reducible structure is sought, i.e. $A,B$ are vertex subsets of $V$. It is crucial that $A$ already separates $B$ from $V\setminus B$. Then, the vertex set $A_{1}\subseteq A$ (cf. Definition 1) typically represents the structure targeted for reduction and can be seen as the “head” of a crown decomposition. The components within $\mathbb{CC}(G[B])$ assigned to $A_{1}$ (the separated part through $A_{1}$), denoted by $f^{-1}(A_{1})$, form the “crown”. In the context of optimization problems such as the vertex cover problem, a successful reduction often involves the head being part of an optimal solution. Subsequently, upon its removal, the crown — separated by the head — no longer needs to be considered.
Ideally, we wish to have $A_{1}=A$, but even when this is not the case the balanced part of a DBE ensures that the elements mapped to $A_{2}$ are bounded, which finally allows us to bound the number of vertices (or the sum of the vertex weights) of $A_{2}\cup V(f^{-1}(A_{2}))$.
Consider a crown decomposition, which is a partition of the vertex set into body, head and crown, where the head separates the body from the crown. The head and crown of the classical crown decomposition corresponds to a $(1,1)$-CD, the $q$-expansion for $q\in\mathds{N}$ to a $(q,1)$-CD and the weighted crown decomposition for $q\in\mathds{N}$ to a $(q,x)$-CD. The last known structure captured by Definition 1 is the balanced expansion which corresponds for $x,y\in\mathds{N}$ to a $(A_{1},A_{2},y,f,x)$ DBE. The essential new structural property of a DBE compared to a balanced expansion are the varying demands for $A$.
Let $G=(A\cup B,E)$ be a graph. For $A^{\prime}\subseteq A$ we define $\mathcal{B}_{A^{\prime}}$ as the components $Q\in\mathbb{CC}(G[B])$ with $N(Q)\subseteq A^{\prime}$. The following theorem, easily derived as modification of the results of [DBLP:conf/esa/Casel0INZ21], gives the runtime to find a DBE and the existence guarantee of $A_{1}$ depending on the size or weight in the graph. For the sake of completeness, we give the proof in Appendix LABEL:app:a.
Theorem 1 (Demanded balanced expansion).
Let $G=\left(A\cup B,E,w\right)$ be a graph with no isolated components in $\mathbb{CC}(G[B])$, i.e. every component of $\mathbb{CC}(G[B])$ contains at least one neighbor of $A$. Let $y\geq\max_{Q\in\mathbb{CC}(G[B])}w(Q)$ and $D=\{d_{a}\}_{a\in A}$ demands with $d_{a}\in\mathds{N}$ for each $a\in A$. A demanded balanced expansion ($A_{1},A_{2},y,f,D$) can be computed in $\mathcal{O}\left(|V|\,|E|\right)$ time. Furthermore, if there is an $A^{\prime}\subseteq A$ with $w(A^{\prime})+w(V(\mathcal{B}_{A^{\prime}}))\geq\sum_{a\in A^{\prime}}d_{a}$, then $A_{1}\neq\varnothing$.
The next graph structure that we use for our kernelization algorithms is introduced by Casel et al. [DBLP:conf/esa/Casel0INZ21] and is a combination of a balanced connected partition and a weighted crown decomposition, which is called balanced crown decomposition. Formally it is defined as follows (see also Figure 1 (right) for an illustration).
Definition 2 ($\lambda$-balanced crown decomposition).
A $\lambda$-balanced crown decomposition of a graph $G=(V,E,w)$ is a tuple $(C,H,\mathfrak{R},f)$, where $\{H,C,R\}$ is a partition of $V$, the set $\mathfrak{R}$ is a partition of $R$, and $f\colon\mathbb{CC}(G[C])\rightarrow H$, such that:
  • 1.
    w(Q)λ for each QCC(G[C]),
  • 2.
    f(Q)N(Q) for each QCC(G[C]),
  • 3.
    N(C)H,
  • 4.
    >+w(h)w(f-1(h))λ for each hH and
[…]
($\lambda,\lambda$)-CD
  • 5.
    $G[R^{\prime}]$ is connected and $\lambda<w(R^{\prime})\leq 3\lambda$ for each $R^{\prime}\in\mathfrak{R}$.
We use $\lambda$-BCD as an abbreviation for a $\lambda$-balanced crown decomposition. Looking at the original definition of a $\lambda$-BCD in [DBLP:conf/esa/Casel0INZ21] (Definition 6), we shift the $\lambda$ value by one, which allows us to change several inequalities between strict and non-strict for a clearer representation, while still keeping the definition the same. Furthermore, $\lambda$ must be at least two in the original definition, but to simplify the understanding of the application to the problems considered in this paper, it makes more sense that a $\lambda$-BCD can also exist with $\lambda=1$.
The authors in [DBLP:conf/esa/Casel0INZ21] provide an algorithm that finds a $\lambda$-BCD in polynomial time, as given in the following.
Theorem 2 (Balanced crown decomposition theorem, [DBLP:conf/esa/Casel0INZ21] Theorem 7).
Let $G=(V,E,w)$ be a graph and $\lambda\in\mathds{N}$, such that each connected component in $G$ has weight larger than $\lambda$. A $\lambda$-balanced crown decomposition $(C,H,\mathfrak{R},f)$ of $G$ can be computed in $\mathcal{O}\left(r^{2}\,|V|\,|E|\right)$ time, where $r=|H|+|\mathfrak{R}|<|V|$ is at most the size of a maximum $(\lambda+1)$-packing.
We end the preliminary section with a formal definition of the subgraph packing problem. Given a graph $G=(V,E)$ and two parameters $r,\lambda\in\mathds{N}$. We say that $P_{1},\dots,P_{m}\subseteq V$ is a $\lambda$-packing if for all $i,j\in[m]$ with $i\neq j$ the induced subgraph $G[P_{i}]$ is connected, $|P_{i}|\geq\lambda$, and $P_{i}\cap P_{j}=\varnothing$. The task is to find a $\lambda$-packing of size at least $r$.
3 Improved Kernels for VI, wVI and wCOC
The vertex integrity of a graph models finding an optimal balance between removing vertices and keeping small connected parts of a graph. As a reminder: In the formal definition, a graph $G$ and a parameter $p$ are given. The task is to find a vertex set $S\subseteq V$ such that $|S|+\max_{Q\in\mathbb{CC}(G-S)}|Q|$ is at most $p$. In the weighted vertex integrity problem (wVI), i.e. the vertices have weights, the aim is that $w(S)+\max_{Q\in\mathbb{CC}(G-S)}w(Q)$ is at most $p$. We improve the vertex kernel of $p^{3}$ provided by Drange et al. [DBLP:journals/algorithmica/DrangeDH16] for both variants VI and wVI. To obtain such a kernel, they essentially established that vertices $v\in V$ with $w(v)+w(N(v))>p$ belong to a solution if we have a yes-instance in hand. A simple counting argument after removing those vertices provides a $p^{3}$ vertex kernel. We improve this by employing crown decompositions, which offer valuable insights into the structural properties of vertex sets rather than focusing solely on individual vertices. Directly applying this method to problems like COC, as seen in prior works such as [DBLP:conf/esa/Casel0INZ21, DBLP:journals/tcs/ChenFSWY19, DBLP:conf/iwpec/KumarL16, DBLP:conf/tamc/XiaoK17], is challenging because we lack prior knowledge about the size of remaining components after solution removal. However, if we had at least a lower bound, we could theoretically safely reduce our instance accordingly. To establish such a bound, we engage in an interplay between packings and separators, where the balanced crown decomposition (BCD) proves instrumental. By identifying reducible structures in the input instance through embedding a DBE into BCD after determining a suitable bound, we prove the following theorems regarding VI and wVI.
Theorem 3.
The vertex integrity problem admits a vertex kernel of size $3p^{2}$ in time $\mathcal{O}\left(\log(p)|V|^{4}|E|\right)$.
Theorem 4.
The weighted vertex integrity problem admits a vertex kernel of size $3(p^{2}+p^{1.5}p_{\ell})$ in time $\mathcal{O}\left(\log(p)|V|^{4}|E|\right)$, where $p_{\ell}$ is at most the size of the largest component after removing a solution.
Closely related to wVI is the weighted component order connectivity problem (wCOC). Given a vertex-weighted graph $G=(V,E,w)$ and two parameters $k,W\in\mathds{N}$, the task is to find a vertex set $S\subseteq V$ such that $w(S)\leq k$ where each component weighs at most $W$. The techniques employed to derive the kernel for wVI can be seamlessly applied to wCOC, thereby enhancing the current state of the art vertex kernel of $kW(k+W)+k$. This kernel is also provided by Drange et al. [DBLP:journals/algorithmica/DrangeDH16] in a similar way as for wVI.
Theorem 5.
The weighted component order connectivity problem admits a vertex kernel of size $3\mu(k+\sqrt{\mu}W)$, where $\mu=\max(k,W)$. Furthermore, such a kernel can be computed in time $\mathcal{O}\left(r^{2}k|V||E|\right)$, where $r$ is the size of a maximum $(W+1)$-packing.
Before we prove these theorems, we give some notations that we use in this section. We define them for the weighted case, as the unweighted case can be viewed in the same way with unit weights. We say that $S$ is a solution if it satisfies $w(S)+\max_{Q\in\mathbb{CC}(G-S)}w(Q)\leq p$ for wVI or $w(S)\leq k$ and $\max_{Q\in\mathbb{CC}(G-S)}w(Q)\leq W$ for wCOC. We denote an instance of wVI by $(G,p)$ and an instance of wCOC by $(G,k,W)$. We say that $(G,p)$ or $(G,k,W)$ is a yes-instance if there is a solution, otherwise, we say that it is a no-instance. Let $\mathcal{S}\subseteq 2^{V(G)}$ be all solutions for $(G,p)$. We define $p_{\ell}:=\min_{S\in\mathcal{S}}\left(\max_{Q\in\mathbb{CC}(G-S)}w(Q)\right)$ which is the optimum lower bound on the size of the connected components after the removal of any solution; where for no-instances we set $p_{\ell}=p$. For all instances $(G,p)$ of wVI, we assume $w(v)<p$ for every $v\in V$ and that $G$ contains a connected component of weight more than $p$. Lastly, for a better readability, we will consistently assume that when computing a $\lambda$-BCD, we disregard small components, meaning that if $(C,H,\mathfrak{R},f)$ is a $\lambda$-BCD for $G$, then $C\cup H\cup V(\mathfrak{R})$ are only the vertices that are contained in some component of size (or weight) more than $\lambda$ of $G$.
3.1 Vertex Integrity
The rest of this section is dedicated to the proof of Theorem 3. We first give the types of reductions we want to use, always with respect to some $c\leq p_{\ell}$, and then discuss how to efficiently compute an estimate of $c$ to apply these reduction rules.
Lemma 6.
Let $(G,p)$ be an instance of VI where not all connected components are of size at most $p$ and let $(C,H,f)$ be a $(p,c)$-CD in $G$ for some $c\leq p_{\ell}$, such that $N(C)\subseteq H$. Then, $(G,p)$ is a yes-instance if and only if $(G-(H\cup C)+K_{c},p-|H|)$ is a yes-instance.
Proof 3.7.
Let $G^{\prime}=G-\{H\cup C\}+K_{c}$. First assume $(G,p)$ is a yes-instance, and let $S$ be an according solution with $s=\max_{Q\in\mathbb{CC}(G-S)}|Q|$. Note that since not all components of $G$ have size at most $p$, it follows that $S\neq\emptyset$ and $s<p$. We claim that $S^{\prime}=S\setminus(H\cup C)$ is a solution for $(G^{\prime},p-|H|)$. It suffices to show that $\max_{Q\in\mathbb{CC}(G^{\prime}-S^{\prime})}|Q|\leq s$ and $|S^{\prime}|\leq|S|-|H|$. More precisely, we can then conclude that $S^{\prime}$ has size at most $|S^{\prime}|\leq(p-|H|)-s$ and creates components of size at most $s$ when deleted from $G^{\prime}$, which makes it a solution for $(G^{\prime},p-|H|)$, meaning that $(G^{\prime},p-|H|)$ is a yes-instance.
For the claim $\max_{Q\in\mathbb{CC}(G^{\prime}-S^{\prime})}|Q|\leq s$, first note that $c\leq p_{\ell}$ by the choice of $c$, and that by the definition of $p_{\ell}$ we have $p_{\ell}\leq s$. Thus the $K_{c}$ component of $G^{\prime}$ has size at most $s$. For the other components in $G^{\prime}-S^{\prime}$, suppose that $G-(H\cup C)-S^{\prime}$ contains a connected component $Q^{\prime}$ of size larger than $s$. Note that $V(Q^{\prime})\subset V\setminus(H\cup C)$. Since we do not remove vertices from $S$ in $V\setminus(C\cup H)$ to build $S^{\prime}$, we have that $Q^{\prime}$ is also a component of $G-S$, contradicting $s=\max_{Q\in\mathbb{CC}(G-S)}|Q|$.
For the claim $|S^{\prime}|\leq|S|-|H|$ observe that $\mathcal{V}_{H}=\{\{h\}\cup V(f^{-1}(h))\}_{h\in H}$ forms a $p$-packing of size $|H|$ in $G[C\cup H]$. Since $p>s$ and the elements in $\mathcal{V}_{H}$ are pairwise disjoint, every element of $\mathcal{V}_{H}$ contains at least one vertex of $S$. Thus, we obtain $|S\cap(C\cup H)|\geq|H|$ and therefore $|S^{\prime}|\leq|S|-|H|$.
For the other direction, assume $(G^{\prime},p-|H|)$ is a yes-instance and let $S^{\prime}$ be a solution for $(G^{\prime},p-|H|)$, and let $\max_{Q\in\mathbb{CC}(G^{\prime}-S^{\prime})}|Q|=s^{\prime}$. We claim that $S=(S^{\prime}\cup H)\cap V$ is a solution for $(G,p)$.
Components of $G-S$ are either the components of $G[C\cup H]-H$, since $H$ separates $C$ from the rest of $G$, or components of $G^{\prime}-S^{\prime}$. By definition of the crown decomposition, all components of $G[C\cup H]-H$ are of size at most $c$, while all components of $G^{\prime}-S^{\prime}$ are of size at most $s^{\prime}$ by the definition of $s^{\prime}$. If $c\leq s^{\prime}$ it follows that all components of $G-S$ have size at most $s^{\prime}\leq p-|H|-|S^{\prime}|\leq p-|S|$ (since $S^{\prime}$ is a solution for $(G^{\prime},p-|H|)$ creating components of size $s^{\prime}$, meaning that $s^{\prime}+|S^{\prime}|\leq p-|H|$, and $|S|\leq|S^{\prime}|+|H|$), and we are done.
Assume that $s^{\prime}<c$. This means that the $K_{c}$ component in $G^{\prime}$ is shrunk to have size at most $s^{\prime}$, meaning that $S^{\prime}$ contains at least $c-s^{\prime}$ vertices from $K_{c}$. This gives $|S|\leq|S^{\prime}|+|H|-c+s^{\prime}$. If $G-S$ now contains a component larger than $s^{\prime}$, this component originates from the crown and has a size of at most $c$. This ensures that $S$ is a solution for $(G,p)$ because
$|S|+c$ $\leq|S^{\prime}|+|H|-c+s^{\prime}+c$
$=|S^{\prime}|+s^{\prime}+|H|$
$\leq p-|H|+|H|$
$\leq p.$
We will eventually use BCDs to find a good estimate for $c$ and to find a CD that can be removed with Lemma 6. Since we only search for $\lambda$-BCDs in components of size larger than $\lambda$, we still need another reduction rule for small components. Clearly, if an instance contains more than $p$ components of size at most $c$, deleting a smallest such component is a safe rule. With the same trick to add a $K_{c}$ like in Lemma 6 we can go further and delete all other small components.
Lemma 3.8.
Let $(G,p)$ be an instance of VI, $c<p_{\ell}$, and let $U$ be a connected component of $G$ of size at most $c$. Then $(G,p)$ is a yes-instance if and only if $(G-U+K_{c},p)$ is a yes-instance.
Proof 3.9.
If $(G,p)$ is a yes-instance, let $S$ be a corresponding solution, meaning that for $s=\max_{Q\in\mathbb{CC}(G-S)}|Q|$, we have $|S|+s\leq p$. By the definition of $p_{\ell}$, we know that $s\geq p_{\ell}$. Each component of $G-U+K_{c}-S$ is a component of $G-S$, and thus has size at most $s$, and the $K_{c}$, which has size $c\leq p_{\ell}\leq s$. Therefore, $S$ is also a solution for $(G-U+K_{c},p)$.
For the other direction, assume $(G-U+K_{c},p)$ is a yes-instance and let $S$ be such that $s=\max_{Q\in\mathbb{CC}(G-U+K_{c}-S)}|Q|$ satisfies $|S|+s\leq p$. We claim that $S^{\prime}=S\cap V(G)$ is a solution for $(G,p)$. Since $U$ is a connected component of $G$, all components in $G-S^{\prime}$, except for $U$, are components of $G-U+K_{c}-S$, so they have size at most $p-|S|\leq p-|S^{\prime}|$. Thus if $|U|\leq s$ then it immediately follows that all components of $G-S^{\prime}$ have size at most $p-|S^{\prime}|$ meaning that $S^{\prime}$ is a solution for $(G,p)$.
If $|U|>s$ then $U$ is a largest component in $G-S^{\prime}$. Since $|U|\leq c$, it follows that $c>s$ which means that the $K_{c}$ component in $G-U+K_{c}$ must be shrunk by $S$. More precisely, $S$ contains at least $c-s$ of its vertices. This means $|S^{\prime}|\leq|S|-c+s$ which gives $|U|+|S^{\prime}|\leq c+|S|-c+s=|S|+s\leq p$. With $U$ being a largest component in $G-S^{\prime}$ it again follows that $S^{\prime}$ is a solution for $(G,p)$.
Note that exhaustive application of Lemma 3.8 creates an equivalent instance with only one component of size at most $c$ (isomorphic to $K_{c}$).
In order to use our two reduction rules, we must first determine a suitable lower bound for $p_{\ell}$. Additionally, we need to ensure the existence of a corresponding weighted crown decomposition when the input graph size exceeds a certain size, and that we can find it efficiently. This is where the BCD comes into play. For $\lambda\in\mathds{N}$ a $\lambda$-BCD $(C,H,\mathfrak{R},f)$ in $G$ can only be computed within the components of $G$ that have a size (or weight) greater than $\lambda$. As a reminder, when we compute a $\lambda$-BCD, we disregard small components, meaning that if $(C,H,\mathfrak{R},f)$ is a $\lambda$-BCD for $G$, then $C\cup H\cup V(\mathfrak{R})$ are only the vertices that are contained in some component of size more than $\lambda$ of $G$.
We specify the following lemma directly in the weighted version to also use it later. The lemma provides a lower bound for $p_{\ell}$ from a $\lambda$-BCD $(C,H,\mathfrak{R},f)$, where we essentially use that a $\lambda$-BCD is also a $(\lambda+1)$-packing of size $|H|+|\mathfrak{R}|$.
Lemma 3.10.
Let $(G,p)$ be an instance of wVI and for $\lambda\in[p]$ let $(C,H,\mathfrak{R},f)$ be a $\lambda$-BCD in $G$. If $|H|+|\mathfrak{R}|>p$, then $\lambda<p_{\ell}$ for the instance $(G[C\cup H\cup V(\mathfrak{R})],p)$.
Proof 3.11.
Let $V^{\prime}=C\cup H\cup\mathfrak{R}^{\prime}$ and consider the graph $G[V^{\prime}]$ (subgraph of $G$ created by removing all components of weight at most $\lambda$). The vertex sets $\mathfrak{R}\cup\{\{h\}\cup V(f^{-1}(h))\}_{h\in H}$ form a $(\lambda+1)$-packing of size $|\mathfrak{R}|+|H|>p$ in $G[V^{\prime}]$. It follows that we need more than $p$ vertices for a vertex separator $S$ in $G[V^{\prime}]$, such that $\max_{Q\in\mathbb{CC}(G-S)}w(Q)\leq\lambda$. Thus for any $U\subset V$ with $|U|\leq p$ it follows that $\max_{Q\in\mathbb{CC}(G-U)}w(Q)>\lambda$. This implies $\lambda<p_{\ell}$ for the instance $(G[V^{\prime}],p)$, since any solution has cardinality at most $p$.
An additional advantage of a $(\lambda+1)$-BCD $(C,H,\mathfrak{R},f)$ is that the balanced part $\mathfrak{R}$ can be upper bounded by $3(\lambda+1)|\mathfrak{R}|\leq 3p_{\ell}|\mathfrak{R}|<3p|\mathfrak{R}|$, while a suitable $\lambda$ choice also upper bounds $|\mathfrak{R}|$ by $p$. In particular, if $H=C=\varnothing$ and $|\mathfrak{R}|\leq p$, then we would already have an instance with a size (or weight) of at most $3p^{2}$.
Clearly, a yes-instance cannot have a lower bound for $p_{\ell}$ larger than $p$ as stated in the following corollary.
Corollary 3.12.
Let $(G,p)$ be an instance of wVI. If for a $p$-BCD $(C,H,\mathfrak{R},f)$ it holds that $|H|+|\mathfrak{R}|>p$, then $(G,p)$ is a no-instance.
The next lemma shows under which conditions we can find a $(p,\lambda)$-CD in $G$ with respect to a $\lambda$-BCD and the current graph size.
Lemma 3.13.
Let $(G,p)$ be an instance of VI and for $\lambda\in[p]$ let $(C,H,\mathfrak{R},f)$ be a $\lambda$-BCD in $G$ with $|H|+|\mathfrak{R}|\leq p$. If $|C|+|H|+|V(\mathfrak{R})|\geq p\max\{3\lambda,p+\lambda\}$, then $G^{\prime}=G[C\cup H]$ contains a $(p,\lambda)$-CD in $G$. Furthermore, we can extract it from $G^{\prime}$ in time $\mathcal{O}(|V(G^{\prime})||E(G^{\prime})|)\subseteq\mathcal{O}(|V(G)||E(G)|)$.
Proof 3.14.
Consider Theorem 1 on $G^{\prime}=G[H\cup C]$ (so $H$ in the role of $A$ and $C$ in the role of $B$), and with $d_{h}=p-1+\lambda$ and $y=\lambda$. This theorem then gives us a DBE $(H_{1},H_{2},\lambda,f,p-1+\lambda)$, with the guarantee $H_{1}\neq\emptyset$ in particular if $|H|+|C|\geq\sum_{h\in H}d_{h}$ (case $A=A^{\prime}$ in the Theorem). The part $(C_{1},H_{1},f)$ with $C_{1}=f^{-1}(H_{1})$ then gives us the desired non-empty $(H_{1},\emptyset,\lambda,f,p-1+\lambda)$ DBE, in other words a $(p,\lambda)$-CD in $G^{\prime}$. Moreover, observe that then $N(C)\subseteq H$ in $G$ and $H_{1}\subseteq H$, with $N(C_{1})\subseteq H_{1}$ in $G^{\prime}$ implies $N(C_{1})\subseteq H_{1}$ in $G$ and hence, $(C_{1},H_{1},f)$ is also a $(p,\lambda)$-CD in $G$.
Thus it remains to show $|H|+|C|\geq\sum_{h\in H}d_{h}=(p-1+\lambda)|H|$. Suppose $|C|+|H|<(p-1+\lambda)|H|$, then:
$|R|+|C|+|H|$ $<3\lambda|\mathfrak{R}|+(p-1+\lambda)|H|$
$\leq 3\lambda|\mathfrak{R}|+(p+\lambda)|H|$
$\leq\max\{3\lambda,p+\lambda\}(|\mathfrak{R}|+|H|)$
$\leq\max\{3\lambda,p+\lambda\}p,$
which is a contradiction to $|R|+|C|+|H|\geq\max\{3\lambda,p+\lambda\}p$.
It follows that $|C|\geq(p-2+\lambda)|H|$ and thus $|H|+|C|\geq|H|+(p-2+\lambda)|H|=\sum_{h\in H}d_{h}$. We can thus use Theorem 1 to extract a $(p,\lambda)$-CD $(C_{1},H_{1},f)$ with $\emptyset\neq H_{1}\subseteq H$ and $C_{1}\subseteq C$ from $G^{\prime}$ in time $\mathcal{O}(|V(G^{\prime})||E(G^{\prime})|)$.
If we now combine Lemma 3.10 and Lemma 3.13, we can finally use Lemma 6 by finding a $\lambda\in[p-1]$ such that a ($\lambda+1$)-BCD $(C,H,\mathfrak{R},f)$ satisfies $|H|+|\mathfrak{R}|\leq p$, while a $\lambda$-BCD $(C^{\prime},H^{\prime},\mathfrak{R}^{\prime},f^{\prime})$ satisfies $|H^{\prime}|+|\mathfrak{R}^{\prime}|>p$, which only adds a factor of $\mathcal{O}(\log(p))$ to the computation cost. Formulated more precisely: For $Q\in\mathbb{CC}(C)$ we have $|Q|\leq\lambda+1\leq p_{\ell}$ as $\lambda<p_{\ell}$ by Lemma 3.10. If the size of $G[C\cup H\cup V(\mathfrak{R})]$ is large, then we can find a ($p,\lambda$)-CD for $G$ in $G[C\cup H]$ in polynomial time by Lemma 3.13, which is a reducible structure by Lemma 6.
With these facts in hand we can design a kernelization algorithm. It takes as input an instance $(G,p)$ of VI and returns an equivalent instance with at most $3p^{2}$ vertices.
Find reducible structures (AlgVI)
  1. 1.
    Compute a $p$-BCD $(C_{0},H_{0},\mathfrak{R}_{0},f_{0})$. If $|H_{0}|+|\mathfrak{R}_{0}|>p$, return a trivial no-instance.
  2. 2.
    Compute a $1$-BCD $(C_{1},H_{1},\mathfrak{R}_{1},f_{1})$. If $|H_{1}|+|\mathfrak{R}_{1}|\leq p$, set $c=1$
  3. 3.
    Otherwise, for $\lambda\in[p-1]$ use binary search to find a ($\lambda+1$)-BCD $(C_{1},H_{1},\mathfrak{R}_{1},f_{1})$ and a $\lambda$-BCD $(C_{2},H_{2},\mathfrak{R}_{2},f_{2})$, such that $|H_{1}|+|\mathfrak{R}_{1}|\leq p$ and $|H_{2}|+|\mathfrak{R}_{2}|>p$. Set $c=\lambda+1$.
  4. 4.
    If $|C_{1}|+|H_{2}|+|V(\mathfrak{R}_{2})|\geq p\max\{3c,p+c\}$, compute a $(p,c)$-CD $(H^{\prime},C^{\prime},f^{\prime})$ in $G[C_{1}\cup H_{1}]$ and return $(G-(H^{\prime}\cup C^{\prime})+K_{c},p-|H^{\prime}|)$.
  5. 5.
    Delete all components of size at most $c$ and add a copy of $K_{c}$ in $G$.
Lemma 3.15.
For any instance $(G,p)$ with $p\geq 2$ and where not all components of $G$ have size at most $p$, exhaustive application of AlgVI produces an equivalent instance with at most $3p^{2}$ vertices.
Proof 3.16.
We first prove that an application of AlgVI is safe. The correctness of step 1 is implied by Corollary 3.12. Assuming $c$ satisfies $c\leq p_{\ell}$, the correctness of steps 5 and 4 follow directly from exhaustive application of Lemma 3.8 and Lemma 6, respectively.
For correctness, it remains to prove that $c$ satisfies $c\leq p_{\ell}$. Since the input $G$ is a graph with at least one component of size more than $p$, the choice $c=1$ in step 2 certainly satisfies $p_{\ell}\geq c$. For step 3 we first have to show that the claimed value for $\lambda$ with a ($\lambda+1$)-BCD $(C_{1},H_{1},\mathfrak{R}_{1},f_{1})$ and a $\lambda$-BCD $(C_{2},H_{2},\mathfrak{R}_{2},f_{2})$, such that $|H_{1}|+|\mathfrak{R}_{1}|\leq p$ and $|H_{2}|+|\mathfrak{R}_{2}|>p$ exists. For a $\lambda$-BCD computation in a binary search for the correct value for $\lambda$, we cannot ensure that $|H|+|\mathfrak{R}|$ increases with decreasing $\lambda$ values, because the sizes of $H$ and $\mathfrak{R}$ are not necessarily monotonic with respect to $\lambda$. (Note, for example, that the elements in $\mathfrak{R}$ have a size range from $\lambda+1$ to $3\lambda$.) For a successful binary search, however, it is sufficient to know that there is a $\lambda$ that satisfies the desired properties for step 3. Since we only enter this step if the extreme cases ($p$-BCD and 1-BCD) in step 1 and 2 hold, there has to exist at least one such $\lambda$ value in-between. By Lemma 3.10, a $\lambda$-BCD $(C_{2},H_{2},\mathfrak{R}_{2},f_{2})$ with $|H_{2}|+|\mathfrak{R}_{2}|>p$ implies $p_{\ell}>\lambda$. Since $p_{\ell}$ is an integer, it follows that $p_{\ell}\geq\lambda+1=c$.
For the size of the resulting instance, first note that a trivial no-instance has constant size. Otherwise, after step 5 the only component of size at most $c$ in $G$ is one copy of $K_{c}$. For the larger components, step 4 reduces further, if the $(p,c)$-CD is non-empty. By Lemma 3.13, a non-empty $(p,c)$-CD can be found if $|C_{1}|+|H_{2}|+|V(\mathfrak{R}_{2})|\geq p\max\{3c,p+c\}$. Thus exhaustive application of AlgVI terminates with a $c$-BCD $(C_{1},H_{1},\mathfrak{R}_{1},f_{1})$ that satisfies $|C_{1}|+|H_{2}|+|V(\mathfrak{R}_{2})|\leq p\max\{3c,p+c\}$ with a $c\leq p_{\ell}$. The union $C_{1}\cup H_{2}\cup V(\mathfrak{R}_{2})$ contains all vertices of the components in $G$ that have size larger than $c$, while after step 5 only one component of size at most $c$ remains. This gives a total of at most $p\max\{3c,p+c\}+c$ vertices in $G$. With $c\leq p_{\ell}<p$, if follows that the reduced graph contains at most $p(3(p-1))+p-1<3p^{2}$ vertices.
Finally, we prove the specified running time of Theorem 3, which completes the proof of Theorem 3. Note that for a yes-instance we have to call the algorithm AlgVI at most $|V|$ times to guarantee the desired kernel.
Lemma 3.17.
Algorithm AlgVI runs in time $\mathcal{O}(\log(p)|V|^{3}|E|)$.
Proof 3.18.
The calculation of a $\lambda$-BCD for $\lambda\in[p]$ runs in time $\mathcal{O}(|V|^{3}|E|)$ by Theorem 2, since a ($\lambda+1$)-packing has size less than $|V|$. The extraction of a $(p,\lambda)$-CD from it runs in time $\mathcal{O}(|V||E|)$ by Theorem 1. Each step consists of the calculation of these constructions with an additional time expenditure of $\mathcal{O}(\log(p))$ in step 3. As a result, the algorithm AlgVI runs in time $\mathcal{O}(\log(p)|V|^{3}|E|)$.
3.2 Weighted Vertex Integrity and Component Order Connectivity
In this section, we shift our focus to the weighted variants, namely wVI and wCOC. While utilizing the packing size of the associated BCD offers a starting point for deriving a lower bound for VI, it proves insufficient for improvements in the weighted setting. Therefore, we integrate the weight of the separator $H$ within a BCD $(C,H,\mathfrak{R},f)$ into our analysis. Additionally, we incorporate two distinct DBE’s into a BCD in case a reduction is not achievable within the respective setting. This approach enables us to establish a tighter lower bound for $p_{\ell}$, estimate the remaining instance size more accurately to obtain the desired kernelization results (cf. Theorems 4 and 5), or identify a no-instance.
We begin by introducing the two DBE structures mentioned in conjunction with a BCD, which serve as preconditions for the upcoming lemmas. Let $G=(V,E,w)$ be a vertex weighted graph, $a,s,\lambda\in\mathds{N}$ and $(C,H,\mathfrak{R},f)$ a $\lambda$-BCD with $\lambda\in[a]$ and $|H|+|\mathfrak{R}|\leq s$, where $\max_{v\in V}w(v)\leq\max(a,s)=:\mu$. Let $D^{Y}=\{d^{Y}_{h}\}_{h\in H}$ and $D^{Z}=\{d^{Z}_{h}\}_{h\in H}$ be demands, where $d^{Y}_{h}=a-2+\lambda\cdot(w(h)+1)$ and $d^{Z}_{h}=w(h)-1+(\sqrt{s}+1)\lambda$ for each $h\in H$. Let $(Y_{1},Y_{2},\lambda,f_{Y},D^{Y})$ and $(Z_{1},Z_{2},\lambda,f_{Z},D^{Z})$ be DBE’s in $G[C\cup H]$ with $Y_{1},Y_{2}$ as well as $Z_{1},Z_{2}$ are partitions of $H$ and $f_{Y}^{-1}(Y_{1}),f_{Y}^{-1}(Y_{2})$ as well as $f_{Z}^{-1}(Z_{1}),f_{Z}^{-1}(Z_{2})$ are partitions of $\mathbb{CC}(G[C])$.
The concept of two DBE’s embedded in a BCD can be understood as follows: With an appropriate choice of parameters $a$, $s$, and $\lambda$, the substructures $Y_{1}$ and $f_{Y}^{-1}(Y_{1})$ within the input graph of a wVI or wCOC instance form a reducible structure. Note that if $Y_{1}\neq\varnothing$ (or $Z_{1}\neq\varnothing$), these sets act as separators, isolating $f_{Y}^{-1}(Y_{1})$ (or $f_{Z}^{-1}(Z_{1})$) from the rest of the graph, as $H$ separates $C$ from the rest. However, if $Y_{1}=\varnothing$, attention shifts to $Z_{1}$. If the weight $w(Z_{1})$ exceeds a certain threshold, we can analyze the heaviest remaining component after removing a separator with bounded weight. If both $Y_{1}=\varnothing$ and $w(Z_{1})$ is bounded, we are able to bound the entire graph, or more precisely, all parts of the underlying $\lambda$-BCD $(C,H,\mathfrak{R},f)$.
Next, we present two crucial lemmas that serve as the foundation for designing the algorithms needed to prove Theorems 4 and 5. The first lemma aims to estimate the instance size.
Lemma 3.19.
Let $w(Z_{1})\leq 2\mu^{1.5}$. If $Y_{1}=\varnothing$, then $w(\mathfrak{R})+w(C\cup H)<3\mu(s+\sqrt{\mu}\lambda)$.
Proof 3.20.
Let $C_{1}=V(f_{Z}^{-1}(Z_{1}))$ and $C_{2}=V(f_{Z}^{-1}(Z_{2}))$. We will use the demands $D^{Y}$ and $D^{Z}$ in the corresponding demanded balanced expansion to upper bound $w(Z_{1}\cup C_{1})$ and $w(Z_{2}\cup C_{2})$, respectively. Note that $C=C_{1}\cup C_{2}$ and $H=Z_{1}\cup Z_{2}$. Moreover, we use the properties of a $\lambda$-BCD to bound the weight of $R=\bigcup_{V^{\prime}\in\mathfrak{R}}V^{\prime}$ by $3|\mathfrak{R}|\lambda$, since $w(V^{\prime})\leq 3\lambda$ for every $V^{\prime}\in\mathfrak{R}$.
First, observe that $f_{Z}^{-1}(Z_{1})\subseteq f_{Y}^{-1}(Z_{1})$ since the components $f_{Z}^{-1}(Z_{1})$ are separated by $Z_{1}$ and must be assigned to $Z_{1}$ in $f_{Y}$. Note that $Y_{2}=H$ since $Y_{1}=\varnothing$, which implies that $Z_{1}\subseteq Y_{2}$. These facts allow us to bound $w(Z_{1}\cup C_{1})$ by the demands $D_{Y}$. As a result, we obtain
$w(Z_{1}\cup C_{1})$ $=\sum_{h\in Z_{1}}\left(w(h)+w(f_{Z}^{-1}(h))\right)$
$\leq\sum_{h\in Z_{1}}\left(w(h)+w(f_{Y}^{-1}(h))\right)$
$\leq\sum_{h\in Z_{1}}\left(d^{Y}_{h}+\lambda-1\right)$
$\leq\sum_{h\in Z_{1}}\left(a-2+\lambda\cdot(w(h)+1)+\lambda-1\right)$
$<\sum_{h\in Z_{1}}\left(a+2\lambda+w(h)\cdot\lambda\right).$
Furthermore, using the demands $D_{Z}$ we can bound $w(Z_{2}\cup C_{2})$, since for each $h\in Z_{2}$ we have
$w(h)+w(f_{Z}^{-1}(h))\leq d_{h}^{Z}+\lambda-1\leq w(h)-1+(\sqrt{s}+1)\lambda+\lambda-1<w(h)+2\lambda+\sqrt{s}\lambda.$
Now a straight forward calculation provides the proof:
$w(\mathfrak{R})+w(C\cup H)$ $\leq 3|\mathfrak{R}|\lambda+w(C_{1}\cup Z_{1})+w(C_{2}\cup Z_{2})$
$<3|\mathfrak{R}|\lambda+\sum_{h\in Z_{1}}(a+2\lambda+w(h)\cdot\lambda)+\sum_{h\in Z_{2}}(w(h)+2\lambda+\sqrt{s}\lambda)$
$=3|\mathfrak{R}|\lambda+|Z_{1}|(a+2\lambda)+|Z_{2}|(\sqrt{s}\lambda+2\lambda)+\lambda\cdot\sum_{h\in Z_{1}}w(h)+\sum_{h\in Z_{2}}w(h)$
$\leq 3|\mathfrak{R}|\lambda+(|Z_{1}|+|Z_{2}|)2\lambda+|Z_{1}|a+|Z_{2}|\sqrt{s}\lambda+w(Z_{1})\cdot\lambda+w(Z_{2})$
$<3|\mathfrak{R}|a+2|H|a+|Z_{1}|a+\mu^{1.5}\lambda+2\mu^{1.5}\lambda+|Z_{2}|\mu$
$\leq 3|\mathfrak{R}|\mu+2|H|\mu+(|Z_{1}|+|Z_{2}|)\mu+3\mu^{1.5}\lambda$
$\leq 3\mu(|\mathfrak{R}|+|H|)+3\mu^{1.5}\lambda$
$\leq 3\mu(s+\sqrt{\mu}\lambda)$
With respect to the weight of $Z_{1}$, the second lemma is relevant to Lemma 3.19 and helps to establish a more precise lower bound for $p_{\ell}$, or to identify instances where no feasible solutions exist.
Lemma 3.21.
If $w(Z_{1})>2\mu^{1.5}$, then there is no separator $S$ such that $w(S)\leq s$ and $\max_{Q\in\mathbb{CC}(G-S)}w(Q)\leq\lambda$.
Proof 3.22.
Let $Z_{1}^{>\sqrt{\mu}}:=\{h\in Z_{1}\mid w(h)>\sqrt{\mu}\}$ and $Z_{1}^{\leq\sqrt{\mu}}:=Z_{1}\setminus Z_{1}^{>\sqrt{\mu}}$. We have $\left|Z_{1}^{\leq\sqrt{\mu}}\right|\leq|H|\leq s\leq\mu$ and therefore $w\left(Z_{1}^{>\sqrt{\mu}}\right)>2\mu^{1.5}-w\left(Z_{1}^{\leq\sqrt{\mu}}\right)\geq 2\mu^{1.5}-\mu\cdot\sqrt{\mu}=\mu^{1.5}$. It follows that $\left|Z_{1}^{>\sqrt{\mu}}\right|>\sqrt{\mu}\geq\sqrt{s}$ as $w(h)\leq\mu$ and $\mu\geq s$.
Let $V_{h}=\{h\}\cup f_{Z}^{-1}(h)$ for each $h\in Z_{1}$. Fix an $h\in H$. We have $w(h)>\sqrt{\mu}\geq\sqrt{s}$ and $w(V_{h})>w(h)-1+(\sqrt{s}+1)\lambda-\lambda+1=w(h)+\sqrt{s}\lambda$, where $G[V_{h}]$ is connected. Furthermore, observe that if we do not remove $h$, we must remove at least $\sqrt{s}$ vertices to guarantee that $G[V_{h}]$ does not contain any component of weight larger than $\lambda$, since removing $h$ decomposes $G[V_{h}]$ into components with weight at most $\lambda$. In other words, for any $V^{\prime}\subset V_{h}\setminus\{h\}$ with $|V^{\prime}|<\sqrt{s}$ we have $\max_{Q\in\mathbb{CC}(G[V_{h}]-V^{\prime})}w(Q)\geq w(V_{h})-\lambda|V^{\prime}|>w(h)+\lambda\sqrt{s}-\lambda\cdot(\sqrt{s}-1)>\lambda$. That is, to ensure that $G[V_{h}]$ contains no component of weight larger than $\lambda$ we either need to remove $h$ from it, or remove at least $\sqrt{s}$ vertices of $V_{h}\setminus\{h\}$. Hence, both options cost us at least $\sqrt{s}$ in $G[V_{h}]$. Since the vertex sets $\{V_{h}\}_{h\in H}$ are pairwise disjoint and are more than $\sqrt{s}$, a separator $S$ ensuring that there is no component of weight larger than $\lambda$ cost us more than $\sqrt{s}\cdot\sqrt{s}=s$ in $G[Z_{1}\cup f_{Z}^{-1}(Z_{1})]\subseteq G$.
With the key lemmas for the wVI and wCOC problems established, we now demonstrate their applications for each problem and design the corresponding algorithms to achieve the desired kernelizations.
Weighted Component Order Connectivity Problem
We begin by considering wCOC and start with a reduction rule based on a specific DBE. We assume that each vertex weighs at most $W$, since any vertex with a larger weight must necessarily be part of the solution. That is, it is a safe reduction rule for an instance $(G,k,W)$ of wCOC to remove such vertices and update $k$ accordingly. Moreover, we assume that every connected component in an instance of wCOC weighs more than $W$. It is not difficult to see that components with lesser weight can be safely removed.
Lemma 3.23.
Let $(G,k,W)$ be an instance of the weighted component order connectivity problem, $(H,H^{\prime},W,f,D)$ a DBE in $G$ with $D=\{d_{h}\}_{h\in H}$, where $d_{h}=W-2+W\cdot(w(h)+1)$ for each $h\in H$ and for $C=V(f^{-1}(H))$ we have $N(C)\subseteq H$. Then, $(G,k,W)$ is a yes-instance if and only if $(G-(H\cup C),k-w(H),W)$ is a yes-instance.
Proof 3.24.
Let $S$ be a solution of $(G,k,W)$, $S^{\prime}=S\setminus(H\cup C)$ and $G^{\prime}=G-(H\cup C)$. We prove that $S^{\prime}$ is a solution for $(G^{\prime},k,W)$, implying that if $(G,k,W)$ is a yes-instance then $(G^{\prime},k,W)$ is a yes-instance. We need to show that $w(S^{\prime})\leq w(S)-w(H)$ and $\max_{Q\in\mathbb{CC}(G^{\prime}-S^{\prime})}w(Q)\leq W$. For the former, let $V_{h}=\{h\}\cup f^{-1}(h)$ for each $h\in H$. Note that $w(V_{h})>d_{h}-W+1=W-1+W\cdot w(h)$, where each $G[V_{h}]$ is connected. Moreover, observe that if we do not remove $h$, we must remove at least $w(h)$ vertices to guarantee that $G[V_{h}]$ does not contain any component of weight larger than $W$, since removing $h$ decomposes $G[V_{h}]$ into components with weight at most $W$. In other words, for any $V^{\prime}\subset V_{h}\setminus\{h\}$ with $|V^{\prime}|<w(h)$ we have $\max_{Q\in\mathbb{CC}(G[V_{h}]-V^{\prime})}w(Q)\geq w(V_{h})-W|V^{\prime}|>W-1+W\cdot w(h)-W\cdot(w(h)-1)\geq W$ as $W\geq 1$. That is, to ensure that $G[V_{h}]$ contains no component of weight $W+1$ we either need to remove $h$ from it, or remove at least $w(h)$ vertices of $V_{h}\setminus\{h\}$. Since the vertex sets $\{V_{h}\}_{h\in H}$ are pairwise disjoint, we have $w(S\cap(C\cup H))\geq\sum_{h\in H}w(h)=w(H)$ and therefore $w(S^{\prime})\leq w(S)-w(H)$. For the latter, suppose that $G^{\prime}-S^{\prime}$ contains a connected component $G^{\prime\prime}$ of weight larger than $W$. Note that $V(G^{\prime\prime})\subset V\setminus(H\cup C)$. Since we do not remove vertices from $S$ in $V\setminus(C\cup H)$, we have $G^{\prime\prime}\subseteq G-S$, contradicting that $S$ is a solution for $(G,k,W)$.
The other direction follows simply because the components of $G[C\cup H]-H$ are of weight at most $W$ while $H$ separates $C$ from the rest of the graph, i.e., $N(C)\subseteq H$. That is, adding $H$ to a solution $S^{\prime\prime}$ in $G^{\prime}$, which costs us additional $w(H)$, ensures that the heaviest component of $G-(S^{\prime\prime}\cup H)$ weighs at most $W$.
Derived from Lemmas 3.19 and 3.21 we specify the corresponding lemmas with the appropriate parameter selection as corollaries to design the kernelization algorithm. We define the BCD and DBE structures properly for wCOC, but to facilitate a direct comparison, the specific parameter choices regarding Lemmas 3.19 and 3.21 would be $a=W$, $s=k$, and $\lambda=W$. Note that $a=\lambda$ satisfies the precondition $\lambda\in[a]$ and $\max_{v\in V(G)}w(v)\leq W$ for an instance $(G,k,W)$ of wCOC the precondition $\max_{v\in V(G)}w(v)\leq\max(a,s)$ of Lemmas 3.19 and 3.21.
Let $(G,k,W)$ be an instance of wCOC, $(C,H,\mathfrak{R},f)$ a $W$-BCD in $G$ with $|H|+|\mathfrak{R}|\leq k$ and $\mu=\max(k,W)$. Let $D^{Y}=\{d^{Y}_{h}\}_{h\in H}$ and $D^{Z}=\{d^{Z}_{h}\}_{h\in H}$ be demands, where $d^{Y}_{h}=W-2+W\cdot(w(h)+1)$ and $d^{Z}_{h}=w(h)-1+(\sqrt{k}+1)W$ for each $h\in H$. Let $(Y_{1},Y_{2},W,f_{Y},D^{Y})$ and $(Z_{1},Z_{2},W,f_{Z},D^{Z})$ be DBE’s in $G[C\cup H]$ with $Y_{1},Y_{2}$ as well as $Z_{1},Z_{2}$ are partitions of $H$ and $f_{Y}^{-1}(Y_{1}),f_{Y}^{-1}(Y_{2})$ as well as $f_{Z}^{-1}(Z_{1}),f_{Z}^{-1}(Z_{2})$ are partitions of $\mathbb{CC}(G[C])$. Note that $(Y_{1},Y_{2},W,f_{Y},D^{Y})$ coincide with the DBE in Lemma 3.29. As a reminder, we assumed that only vertices in components of $G$ with weight greater than $W$ are part of the $W$-BCD. Since we have already established that $G$ contains only components with weight greater than $W$, it follows that all vertices of $G$ are part of a $W$-BCD in $G$.
Corollary 3.25.
Let $(G,k,W)$ be an instance of wCOC and let $w(Z_{1})\leq 2\mu^{1.5}$. If $w(V(G))\geq 3\mu(k+\sqrt{\mu}W)$, then $Y_{1}\neq\varnothing$.
Corollary 3.26.
Let $(G,k,W)$ be an instance of wCOC. If $w(Z_{1})>2\mu^{1.5}$, then $(G,k,W)$ is a no-instance.
The algorithm takes as input an instance $(G,k,W)$ of wCOC, where $w(G)\geq 3\mu(k+\sqrt{\mu}W)$. It either outputs a reduced instance or determines that $(G,k,W)$ is a no-instance by returning a trivial no-instance. Note that the weight condition on $w(G)$ is not a loss of generality, as there is nothing to do if $w(G)<3\mu(k+\sqrt{\mu}W)$.
Find reducible structures (AlgWCOC):
  1. 1.
    Compute a $W$-BCD $(C,H,\mathfrak{R},f)$ in $G$. If $|H|+|\mathfrak{R}|>k$, then return a trivial no-instance.
  2. 2.
    Let $d_{h}=W-2+W\cdot(w(h)+1)$ for each $h\in H$ and $D=\{d_{h}\}_{h\in H}$. Compute a DBE $(H_{1},H_{2},W,f,D)$ in $G[C\cup H]$ with $H_{1},H_{2}\subseteq H$ and $f^{-1}(H_{1}),f^{-1}(H_{2})\subseteq C$.
  3. 3.
    If $H_{1}=\varnothing$, then return a trivial no-instance. Otherwise, return $(G-(H_{1}\cup f^{-1}(H_{1})),k-w(H_{1}),W)$.
With the following lemma we prove Theorem 5. Note that we need to apply AlgWCOC at most $k$ times.
Lemma 3.27.
Algorithm AlgWCOC works correctly and has a runtime of $\mathcal{O}(r^{2}|V||E|)$, where $r$ is the size of a maximum $(W+1)$-packing.
Proof 3.28.
If $|H|+|\mathfrak{R}|>k$, then the algorithm returns a trivial no-instance. This is correct as $\mathfrak{R}\cup\{\{h\}\cup V(f^{-1})(h)\}_{h\in H}$ forms a $(W+1)$-packing in $G$ of size larger than $k$, which in turn shows that we need to remove more than $k$ vertices to achieve that the remaining graph contains only components of weight $W$.
If $H_{1}=\varnothing$, then the case in Corollary 3.26 must be true, as otherwise, by the weight of $G$ and Corollary 3.25 we must have $H_{1}\neq\varnothing$ (in the corollary $Y_{1}\neq\varnothing$). This shows that the conclusion that $(G,k,W)$ is a no-instance for the case $H_{1}=\varnothing$ is also correct.
Otherwise, the algorithm reduces the instance $(G,k,W)$ to $(G-(H_{1}\cup f^{-1}(H_{1})),k-w(H_{1}),W)$, where the reduction is safe by Lemma 3.23.
Regarding the running time, computing a BCD or a DBE as well as all other steps take not more time than $\mathcal{O}(r^{2}|V||E|)$ (cf. Theorems 1 and 2).
Weighted Vertex Integrity Problem
We now turn our attention to wVI and prove Theorem 4. As a reminder, we assume for an instance $(G,p)$ of wVI that there is a component of weight greater than $p$ and that every vertex has weight lesser than $p$; otherwise, it can be immediately labeled as a trivial yes-instance or no-instance, respectively. Again, we begin with a reduction rule based on a specific DBE. We point out that one direction of the reduction is very similar to the reduction in Lemma 3.23 for wCOC. However, the other direction requires more careful consideration.
Lemma 3.29.
Let $(G,p)$ be an instance of the weighted vertex integrity problem, $\lambda\in\mathds{N}$, where $1\leq\lambda\leq p_{\ell}$, and $(H,H^{\prime},\lambda,f,D)$ a DBE in $G$ with $D=\{d_{h}\}_{h\in H}$, where $d_{h}=p-2+\lambda\cdot(w(h)+1)$ for each $h\in H$ and for $C=V(f^{-1}(H))$ we have $N(C)\subseteq H$. Then, $(G,p)$ is a yes-instance if and only if $(G-(H\cup C)+v,p-w(H))$ is a yes-instance, where $v$ is a vertex of weight $\lambda$.
Proof 3.30.
Let $S$ be a solution of $(G,p)$ with $c=\max_{Q\in\mathbb{CC}(G-S)}w(Q)$, $S^{\prime}=S\setminus(H\cup C)$ and $G^{\prime}=G-(H\cup C)+v$. We prove that $S^{\prime}$ is a solution for $(G^{\prime},p-w(H))$ with $\max_{Q\in\mathbb{CC}(G^{\prime}-S^{\prime})}w(Q)\leq c$, implying that if $(G,p)$ is a yes-instance then $(G^{\prime},p-w(H))$ is a yes-instance. We need to show that $w(S^{\prime})\leq w(S)-w(H)$ and $\max_{Q\in\mathbb{CC}(G^{\prime}-S^{\prime})}w(Q)\leq c$. For the former, let $V_{h}=\{h\}\cup f^{-1}(h)$ for each $h\in H$. Note that $w(V_{h})>d_{h}-\lambda+1=p-1+\lambda\cdot w(h)$, where each $G[V_{h}]$ is connected. Moreover, observe that if we do not remove $h$, we must remove at least $w(h)$ vertices to guarantee that $G[V_{h}]$ does not contain any component of weight larger than $p$, since removing $h$ decomposes $G[V_{h}]$ into components with weight at most $\lambda$. In other words, for any $V^{\prime}\subset V_{h}\setminus\{h\}$ with $|V^{\prime}|<w(h)$ we have $\max_{Q\in\mathbb{CC}(G[V_{h}]-V^{\prime})}w(Q)\geq w(V_{h})-\lambda|V^{\prime}|>p-1+\lambda\cdot w(h)-\lambda\cdot(w(h)-1)\geq p$ as $\lambda\geq 1$. That is, to ensure that $G[V_{h}]$ contains no component of weight $p+1>c$ we either need to remove $h$ from it, or remove at least $w(h)$ vertices of $V_{h}\setminus\{h\}$. Since the vertex sets $\{V_{h}\}_{h\in H}$ are pairwise disjoint, we have $w(S\cap(C\cup H))\geq\sum_{h\in H}w(h)=w(H)$ and therefore $w(S^{\prime})\leq w(S)-w(H)$. For the latter, suppose that $G^{\prime}-S^{\prime}$ contains a connected component $G^{\prime\prime}$ of weight larger than $c$. Note that $V(G^{\prime\prime})\subset V\setminus(H\cup C)$. Moreover, note that the added vertex $v$ in $G^{\prime}$ is isolated and of weight $\lambda\leq p_{\ell}\leq c$ and thus, $v\notin V(G^{\prime\prime})$. Since we do not remove vertices from $S$ in $V\setminus(C\cup H)$, we have $G^{\prime\prime}\subseteq G-S$, contradicting $c=\max_{Q\in\mathbb{CC}(G-S)}w(Q)$.
For the other direction, we first show that if $(G^{\prime},p-w(H))$ is a yes-instance, then there is always a solution $R$ such that $\max_{Q\in\mathbb{CC}(G^{\prime}-R)}w(Q)\geq\lambda$. Once we have this, the other direction follows simply because the components of $G[C\cup H]-H$ have weight at most $\lambda$, while $H$ separates $C$ from the rest of the graph, i.e., $N(C)\subseteq H$. Adding $H$ to $S^{\prime\prime}$ incurs an additional cost of $w(H)$, which shows that $R\cup H$ is a solution for $(G,p)$.
Let $R^{\prime}$ be a solution of $(G^{\prime},p-w(H))$ with $c^{\prime}=\max_{Q\in\mathbb{CC}(G^{\prime}-R^{\prime})}w(Q)$. If $c^{\prime}\geq\lambda$, then there is nothing to do, and we are done. Otherwise, $v\in R^{\prime}$, since $v$ is isolated in $G^{\prime}$ and has weight $\lambda$. We remove $v$ from $R^{\prime}$, and let the new set be $R^{\prime\prime}$. Note that ($(\{v\},\varnothing)$ is now the heaviest connected component in $G^{\prime}-R^{\prime\prime}$, and therefore $\max_{Q\in\mathbb{CC}(G^{\prime}-R^{\prime\prime})}w(Q)=w(v)=\lambda$. Finally, we need to show that $R^{\prime\prime}$ is a feasible solution in $G^{\prime}$, which is demonstrated by the following computation:
$p-w(H)\geq w(R^{\prime})+c^{\prime}=w(R^{\prime})-\lambda+\lambda+c^{\prime}\geq w(R^{\prime})-w(v)+\lambda=w(R^{\prime\prime})+\lambda.$
To design the upcoming algorithm, we need to establish an additional reduction rule, ensuring that isolated components with a weight of at most $p_{\ell}$ can be safely removed.
Lemma 3.31.
Let $(G,p)$ be an instance of wVI that contains at least one component of weight greater than $p$, where $p>1$. Let $c\leq p_{\ell}$ and let $U\subset V(G)$ be all vertices within connected components of weight at most $c$.
  1. 1.
    If $c>1$, then $(G,p)$ is a yes-instance if and only if $(G-U+v,p)$ is a yes-instance, where $w(v)=c$.
  2. 2.
    If $c=1$, then $(G,p)$ is a yes-instance if and only if $(G-U,p)$ is a yes-instance.
Proof 3.32.
  1. 1.
    Let $G^{\prime}=G-U+v$. Let $S$ be a solution of $(G,p)$, and define $x=\max_{Q\in\mathbb{CC}(V(G)-S)}w(Q)$. For $S^{\prime}=S\setminus U$, it follows that $x\geq\max_{Q\in\mathbb{CC}(V(G)-S^{\prime})}w(Q)$, since the largest component of $\mathbb{CC}(G[U])$ has weight $c\leq p_{\ell}\leq x$. As a result, $S^{\prime}$ is a solution for $(G^{\prime},p)$ because $G-U=G^{\prime}-\{v\}$ and $w(v)=c\leq x$.
    Let $R$ be a solution of $(G^{\prime},p)$, let $y=\max_{Q\in\mathbb{CC}(G^{\prime}-R)}w(Q)$, and set $R^{\prime}=R\setminus\{v\}$. First, note that $\max_{Q\in\mathbb{CC}(G^{\prime}-R^{\prime})}w(Q)\leq y$ since $v$ is an isolated vertex in $G^{\prime}$. If $y\geq c$, then $R^{\prime}$ is a solution for $(G,p)$, since $G-U=G^{\prime}-\{v\}$ and $\max_{Q\in\mathbb{CC}(G[U])}w(Q)\leq c\leq y$. Otherwise, $v\in R$ as $y<c$. Now, note that ($(\{v\},\varnothing)$ is the heaviest connected component in $G^{\prime}-R^{\prime}$. It follows that $c=w(v)=\max_{Q\in\mathbb{CC}(G^{\prime}-R^{\prime})}w(Q)\geq\max_{Q\in\mathbb{CC}(G-R^{\prime})}w(Q)$, since $\max_{Q\in\mathbb{CC}(G-U-R^{\prime})}w(Q)\leq y<c$ as $G-U=G^{\prime}-\{v\}$ and $\max_{Q\in\mathbb{CC}(G[U])}w(Q)\leq c$. Finally, we show that $R^{\prime}$ is a feasible solution for $G$, proving the lemma:
    $p\geq w(R)+y=w(R)+y+c-c\geq w(R)-c+c=w(R)-w(v)+c=w(R^{\prime})+c$
  2. 2.
    Let $S$ be a solution of $(G,p)$ and let $G^{\prime}=G-U$. Since there is a component with weight greater than $p$, we have $\max_{Q\in\mathbb{CC}(G-S)}w(Q)\geq 1$. It follows that $S^{\prime}=S\setminus U$ is also a solution for $(G,p)$, since $w(u)=1$ for the isolated vertices $u\in U$. Finally, $S^{\prime}$ is also a solution for $(G^{\prime},p)$ as $G^{\prime}\subseteq G$.
    Let $R$ be a solution of $(G^{\prime},p)$. The graph $G$ contains a component $T$ with weight greater than $p\geq 1$, so $V(T)\cap U=\varnothing$. Since $T$ is also in $G^{\prime}$, we have $\max_{Q\in V(G^{\prime})}w(Q)\geq 1$ because not all vertices of $V(T)$ can be in $R$. As a result, $R$ is also a solution for $(G,p)$, since $w(u)=1$ for the isolated vertices $u\in U$.
Derived from Lemmas 3.19 and 3.21 we specify the corresponding lemmas with the appropriate parameter selection as corollaries to design the kernelization algorithm. We define the BCD and DBE structures properly for wVI, but to facilitate a direct comparison, the specific parameter choices regarding Lemma 3.19 and Lemma 3.21 would be $a=s=p$, and $\lambda$ is not fixed, as we do not know $p_{\ell}$. However, as in the preconditions of Lemmas 3.19 and 3.21, it is assumed that $\lambda\in[a]$. Note also that $\max_{v\in V(G)}w(v)\leq p$ for an instance $(G,p)$ of wVI satisfies the precondition $\max_{v\in V(G)}w(v)\leq\max(a,s)$.
Let $(G,p)$ be an instance of wVI and $(C,H,\mathfrak{R},f)$ a $\lambda$-BCD in $G$ with $|H|+|\mathfrak{R}|\leq p$. Let $D^{Y}=\{d^{Y}_{h}\}_{h\in H}$ and $D^{Z}=\{d^{Z}_{h}\}_{h\in H}$ be demands, where $d^{Y}_{h}=p-2+\lambda\cdot(w(h)+1)$ and $d^{Z}_{h}=w(h)-1+(\sqrt{p}+1)\lambda$ for each $h\in H$. Let $(Y_{1},Y_{2},\lambda,f_{Y},D^{Y})$ and $(Z_{1},Z_{2},\lambda,f_{Z},D^{Z})$ be DBE’s in $G[C\cup H]$ with $Y_{1},Y_{2}$ as well as $Z_{1},Z_{2}$ are partitions of $H$ and $f_{Y}^{-1}(Y_{1}),f_{Y}^{-1}(Y_{2})$ as well as $f_{Z}^{-1}(Z_{1}),f_{Z}^{-1}(Z_{2})$ are partitions of $\mathbb{CC}(G[C])$. Note that $(Y_{1},Y_{2},\lambda,f_{Y},D^{Y})$ coincides with the DBE in Lemma 3.29 only if $\lambda\leq p_{\ell}$.
Corollary 3.33.
Let $(G,p)$ be an instance of wVI and let $w(Z_{1})\leq 2p^{1.5}$. If $w(\mathfrak{R})+w(C)+w(H)\geq 3p(p+\sqrt{p}\lambda)$, then $Y_{1}\neq\varnothing$.
Corollary 3.34.
Let $(G,p)$ be an instance of wVI and let $w(Z_{1})>2p^{1.5}$.
  1. 1.
    If $\lambda<p$, then $\lambda<p_{\ell}$.
  2. 2.
    If $\lambda\geq p$, then $(G,p)$ is a no-instance.
With these corollaries and Lemma 3.10, we can design the following kernelization algorithm. It takes as input an instance $(G,p)$ of wVI, where $|V(G)|>3p(p+\sqrt{p})$ and there is a connected component in $G$ with weight greater than $p$. It either determines that $(G,p)$ is a no-instance by returning a trivial no-instance, or that $(G,p)$ already satisfies the desired kernel, or it outputs a reduced instance. Note that the target vertex kernel contains a factor $p_{\ell}<p$ that we do not know beforehand, which is why an instance may already satisfy the vertex kernel. Moreover, a vertex kernel of $3p(p+\sqrt{p})$ is the smallest size we can guarantee from Theorem 4, as $p_{\ell}\geq 1$. Finally, if there is no component with weight greater than $p$, there is nothing to do. These facts show that the weight assumptions on $G$ are not a loss of generality.
To enhance the readability of the algorithm, we define $\lambda$-crown structures for a $\lambda\in\mathbb{N}$ and demands $D$ to denote a $\lambda$-BCD $(C,H,\mathfrak{R},f)$ and a DBE $(H_{1},H_{2},\lambda,f,D)$ in $G[C\cup H]$, where $H_{1},H_{2}$ forms a partition of $H$ and $f^{-1}(H_{1}),f^{-1}(H_{2})$ a partition of $\mathbb{CC}(G[C])$. As a reminder, when we compute a $\lambda$-BCD, we disregard small components, meaning that if $(C,H,\mathfrak{R},f)$ is a $\lambda$-BCD for $G$, then $C\cup H\cup V(\mathfrak{R})$ are only the vertices that are contained in some component of weight more than $\lambda$ of $G$.
Find reducible structures (AlgWVI):
  1. 1.
    Let $(C^{1},H^{1},\mathfrak{R}^{1},f_{1})$ and $(H^{1}_{1},H^{1}_{2},1,g_{1},\{d^{1}_{h}\}_{h\in H^{1}})$ be $1$-crown-structures with $d^{1}_{h}=w(h)+\sqrt{p}$.
    1. (a)
      If $U=V(G)\setminus(C^{1}\cup H^{1}\cup V(\mathfrak{R}^{1}))$ is not empty, then return $(G-U,p)$.
    2. (b)
      If $w(H^{1}_{1})\leq 2p^{1.5}$ and $|H^{1}|+|\mathfrak{R}^{1}|\leq p$, then compute a DBE $(H^{2}_{1},H^{2}_{2},1,g_{2},\{d^{2}_{h}\}_{h\in H^{1}})$ with $d^{2}_{h}=p-2+w(h)+1$ and return $(G-(H^{2}_{1}\cup g_{2}^{-1}(H^{2}_{1})),p-w(H^{2}_{1}))$.
  2. 2.
    Otherwise, find for $\lambda\in[p-1]$ and for each $h\in H$ with demands
    • $d^{3}_{h}=w(h)-1+(\sqrt{p}+1)\lambda$,
    • $d^{4}_{h}=w(h)-1+(\sqrt{p}+1)(\lambda+1)$
    a $\lambda$-crown-structure and a ($\lambda+1$)-crown-structure
    • $(C^{3},H^{3},\mathfrak{R}^{3},f_{3})$, $(H^{3}_{1},H^{3}_{2},\lambda,g_{3},\{d^{3}_{h}\}_{h\in H})$,
    • $(C^{4},H^{4},\mathfrak{R}^{4},f_{4})$, $(H^{4}_{1},H^{4}_{2},\lambda+1,g_{4},\{d^{4}_{h}\}_{h\in H})$,
    such that
    • $w(H^{3}_{1})>2p^{1.5}$ and/or $|H^{3}|+|\mathfrak{R}^{3}|>p$,
    • $w(H^{4}_{1})\leq 2p^{1.5}$ and $|H^{4}|+|\mathfrak{R}^{4}|\leq p$.
    1. (a)
      If such a $\lambda$ does not exist, then return a trivial no-instance.
    2. (b)
      If $U^{\prime}=V(G)\setminus(C^{4}\cup H^{4}\cup V(\mathfrak{R}^{4}))$ has cardinality larger than one, then return $(G-V^{\prime\prime}+\{v\},p)$, where $v$ is a vertex of weight $\lambda+1$.
    3. (c)
      If $w(V(\mathfrak{R}^{4}))+w(C^{4})+w(H^{4})<3p(p+\sqrt{p}(\lambda+1))$, then report that $|V(G)|\leq 3p(p+\sqrt{p}p_{\ell})$.
  3. 3.
    Otherwise, compute a DBE $(H^{5}_{1},H^{5}_{2},\lambda+1,g_{5},\{d^{5}_{h}\}_{h\in H^{4}})$ with $d^{5}_{h}=p-2+(\lambda+1)\cdot(w(h)+1)$ in $G[C^{4}\cup H^{4}]$ with $H^{5}_{1},H^{5}_{2}\subseteq H^{4}$ and $g_{5}^{-1}(H^{5}_{1}),g_{5}^{-1}(H^{5}_{2})\subseteq C^{4}$ and return $(G-(H^{2}_{1}\cup g_{5}^{-1}(H^{5}_{1})+v),p-w(H^{2}_{1}))$, where $v$ is a vertex of weight $\lambda+1$.
With the following lemma, we prove Theorem 4. Note that we need to apply AlgWVI at most $|V|$ times until it either reports that the input instance already has the desired kernel or, after a reduction, the remaining instance is of size at most $3p(p+\sqrt{p})$. This is because, if the input graph does not have the desired kernel size or is labeled as a no-instance, each application of AlgWVI reduces the input instance by at least one vertex.
Lemma 3.35.
Algorithm AlgWVI works correctly and has a runtime of $\mathcal{O}(\log(p)|V|^{3}|E|)$.
Proof 3.36.
To prove that step 1 works correctly, note that $p_{\ell}\geq 1$ for the instance $(G,p)$, since there is a connected component in $G$ with weight greater than $p$. Step 1a is correct by Lemma 3.31. If in step 1b we have $w(H^{1}_{1})\leq 2p^{1.5}$ and $|H^{1}|+|\mathfrak{R}^{1}|\leq p$, then by Corollary 3.33 and the fact that $w(G)>3p(p+\sqrt{p})$, there must exist a corresponding DBE with $H^{2}_{1}\neq\varnothing$. Additionally, the reduction executed is safe by Lemma 3.29, since $p_{\ell}\geq 1$.
If in step 2 the corresponding $\lambda$ does not exist, then by Corollary 3.12 or Corollary 3.34, $(G,p)$ must be a no-instance, proving that step 2a works correctly. Note that after step 1 we have $w(H^{1}_{1})>2p^{1.5}$ and/or $|H^{1}|+|\mathfrak{R}^{1}|>p$. That is, there is at least for one $\lambda\in[p-1]$ a $\lambda$-BCD $(C,H,\mathfrak{R},f)$ with $w(H)>2p^{1.5}$ and/or $|H|+|\mathfrak{R}|>p$ and if these properties also hold for a ($\lambda+1=p$)-BCD, then by Corollary 3.12 or Corollary 3.34 we have a no-instance as input. Otherwise, a desired $\lambda$ exists and by Lemma 3.10 or Corollary 3.34, we have $\lambda+1\leq p_{\ell}$. All vertices within connected components of weight at least $\lambda+2$ are in $V\left(\mathfrak{R}^{4}\right)\cup C^{4}\cup H^{4}$. The remaining ones, i.e., the vertices within connected components of weight at most $\lambda+1\leq p_{\ell}$, are in $U^{\prime}$. Thus, by Lemma 3.31, step 2b works correctly. Finally, for step 2, we show that step 2c also works correctly. If $w(V(\mathfrak{R}^{4}))+w(C^{4})+w(H^{4})<3p(p+\sqrt{p}(\lambda+1))$, then $|V(G)|\leq 3p(p+\sqrt{p}p_{\ell})$, since $V(\mathfrak{R}^{4})\cup C^{4}\cup H^{4}\cup U^{\prime}=V(G)$, where $|U^{\prime}|\leq 1$ by step 2b.
If we reach step 3, then $w(H^{4}_{1})\leq 2p^{1.5}$, $|H^{4}|+|\mathfrak{R}^{4}|\leq p$ and $w(G)>3p(p+\sqrt{p}(\lambda+1))$. Consequently, by Corollary 3.33, $H^{5}_{1}$ is non-empty, and the executed reduction is safe by Lemma 3.29, since $\lambda+1\leq p_{\ell}$.
The calculation of a $\lambda$-BCD for $\lambda\in[p]$ runs in time $\mathcal{O}(|V|^{3}|E|)$ by Theorem 2, as a $(\lambda+1)$-packing has size less than $|V|$. The extraction of a DBE from it runs in time $\mathcal{O}(|V||E|)$ by Theorem 1. Each step consists of these constructions with an additional time cost of $\mathcal{O}(\log(p))$ in step 2. As a result, the algorithm runs in time $\mathcal{O}(\log(p)|V|^{3}|E|)$.
4 Kernels for Component Order Connectivity
In this section, we consider COC, a refined version of VI. For VI, wVI, and wCOC, crown structures were integrated into the head and crown of the BCD. In contrast, we now incorporate them into the body of the BCD.
Given a graph $G=(V,E)$ and parameters $k,W\in\mathbb{N}$, COC is the problem of removing at most $k$ vertices so that each resulting component has at most $W$ vertices. We present a kernelization algorithm (
Algorithm 1.
2kw
) that produces a $2kW$-vertex kernel in FPT-runtime, parameterized by the size of a maximum $(W+1)$-packing, as stated in the following theorem.
Theorem 4.37.
A vertex kernel of size $2kW$ for the component order connectivity problem can be computed in time $\mathcal{O}(r^{3}|V||E|\cdot r^{\min(3r,k)})$, where $r\leq k$ is the size of a maximum $(W+1)$-packing.
A kernel of size $2kW$ is also presented in [DBLP:conf/iwpec/KumarL16], however with an FPT-runtime in the parameter $W$ and using linear programming methods. In contrast, our result has an FPT-runtime in the parameter of a maximum $(W+1)$-packing $r\leq k$ and is fully combinatorial. Although generalizing our FPT-runtime algorithm to polynomial time remains challenging, our insights into the structural properties of crown decompositions suggest promising directions for future work. Moreover, we demonstrate how our algorithm (AlgCOC-2) becomes polynomial for two cases: when $W=1$ (Vertex Cover) and for claw-free graphs, as stated in the theorems below, whose proofs can be found in Section LABEL:sec:kernels_for_special_COC.
Theorem 4.38.
For $W=1$, i.e. for the vertex cover problem, algorithm AlgCOC-2 provides a $2k$ vertex kernel in polynomial time.
Theorem 4.39.
The component order connectivity problem admits a $2kW$ vertex-kernel on claw-free graphs in polynomial time.
We begin with preliminaries on the reducible structures of COC and their properties, which are essential for our proofs. Following this, we focus on the kernelization algorithms and provide proofs of its correctness.
4.1 Reducible Structures of the Component Order Connectivity Problem and its Properties
We call a (not necessarily minimal) set of vertices that separates a graph into components of size at most $W$ a $W$-separator. We define $(G,k,W)$ as an instance of COC, where it is labeled as a yes-instance if there exists a $W$-separator of size $k$; otherwise, it is labeled as a no-instance. To design the kernelization algorithm, we briefly review, in addition to the preliminary content in Section 2, some definitions and theorems from [DBLP:conf/gecco/Baguley0N0PZ23, DBLP:conf/esa/Casel0INZ21, fomin2019kernelization, DBLP:conf/iwpec/KumarL16]. Of particular interest are the strictly reducible pairs introduced by Kumar and Lokshtanov [DBLP:conf/iwpec/KumarL16], who proved that such structures must exist in graphs with more than $2kW$ vertices.
Definition 4.40 ((strictly) reducible pair).
For a graph $G=(V,E)$ and a parameter $W\in\mathds{N}$, a pair $(A,B)$ of vertex disjoint subsets of $V$ is a reducible pair for COC if the following conditions are satisfied:
  • $N(B)\subseteq A$.
  • The size of each $Q\in\mathbb{CC}(G[B])$ is at most $W$.
  • There is an assignment function $g\colon\mathbb{CC}(G[B])\times A\to\mathds{N}_{0}$, such that
    • for all $Q\in\mathbb{CC}(G[B])$ and $a\in A$, if $g(Q,a)\neq 0$, then $a\in N(Q)$
    • for all $a\in A$ we have $\sum_{Q\in\mathbb{CC}(G[B])}g(Q,a)\geq 2W-1$
    • for all $Q\in\mathbb{CC}(G[B])$ we have $\sum_{a\in A}g(Q,a)\leq|Q|$.
In addition, if there exists an $a\in A$ such that $\sum_{Q\in\mathbb{CC}(G[B])}g(Q,a)\geq 2W$, then $(A,B)$ is a strictly reducible pair.
We say that $(A,B)$ is a minimal (strictly) reducible pair if there does not exist a (strictly) reducible pair $(A^{\prime},B^{\prime})$ with $A^{\prime}\subset A$ and $B^{\prime}\subseteq B$. Clearly, if a graph $G$ contains a (strictly) reducible pair, then it also contains a minimal one. Furthermore, for $A^{\prime}\subseteq A$, we define $\mathbb{CC}(G[B])_{A^{\prime}}\subseteq\mathbb{CC}(G[B])$ as the set of components $\{Q\in\mathbb{CC}(G[B])\mid N(Q)\cap A^{\prime}\neq\varnothing\}$. Before discussing the connection between a (strictly) reducible pair and COC, we state a property of minimal strictly reducible pairs found by Baguley et al. [DBLP:conf/gecco/Baguley0N0PZ23].
Lemma 4.41 ([DBLP:conf/gecco/Baguley0N0PZ23], Lemma 4.16).
Let $(A,B)$ be a minimal strictly reducible pair in $G$ with parameter $W$. Then, for every $A^{\prime}\subseteq A$ we have $|V(\mathbb{CC}(G[B])_{A^{\prime}})|>|A^{\prime}|(2W-1)$.
Considering Lemma 4.41, if there is a minimal strictly reducible pair $(A,B)$ in $G$ and we remove a proper subset $A^{\prime}$ of $A$ from $G$, then $G-A^{\prime}$ still contains a strictly reducible pair with head vertices $A\setminus A^{\prime}$. Specifically, the specified vertex, guaranteed to receive a fractional assignment of at least $2W$ over $g$ from $\mathbb{CC}(G[B])$, can be chosen arbitrarily (cf. Definition 4.40). In our algorithm, these properties enable a stepwise approach in localizing minimal strictly reducible pairs.
A reducible pair $(A,B)$ is, in fact, a $(W,W)$-CD, or equivalently, an $(A,\varnothing,W,f,2W-1)$ DBE where $B=f^{-1}(A)$. Note that $|B|>|A|(2W-1)$; hence, it can be identified in graphs resembling bipartite graphs with the required expansion properties (see Definition 1 and Theorem 1). Here, the head and crown correspond to $A$ and $B$, respectively.
Once a $(W,W)$-CD is identified, we can reduce the COC instance by including the head vertices in the solution. In this case, no additional vertices from the crown are required, since after removing the head vertices, the crown’s connected components are isolated from $G$ and each connected component has size at most $W$.
In the following, we outline why this reduction is safe for COC. Let $P_{1},\dots,P_{m}\subseteq V$ be a $(W+1)$-packing. For a $W$-separator $S\subseteq V$, it holds that $S\cap P_{i}\neq\varnothing$ for all $i\in[m]$. Therefore, the size of a $(W+1)$-packing provides a lower bound on the number of vertices required for a $W$-separator. Considering a $(W,W)$-CD $C,H,f$ in $G$, the function $f$ enables the construction of a $(W+1)$-packing of size $|H|$ in $G[H\cup C]$. This essentially provides a lower bound of $|H|$ vertices for a $W$-separator in $G[H\cup C]$. On the other hand, $H$ serves as a $W$-separator of $G[A\cup B]$ while separating $C$ from the rest of the graph. These properties support the following theorem for $(W,W)$-CDs.
Lemma 4.42 ([DBLP:conf/esa/Casel0INZ21] Lemma 49).
Let $(G,k)$ be an instance of COC and $(A,B,f)$ a $(W,W)$-CD in $G$. Then, $(G,k)$ is a yes-instance if and only if $(G-(A\cup B),k-|A|)$ is a yes-instance.
Finally, the relationship between the existence of a strictly reducible pair and the attainment of the desired kernelization of $2kW$ is demonstrated in the following lemma.
Lemma 4.43 ([fomin2019kernelization], Lemma 6.14).
Let $(G,k,W)$ be an instance of the $W$-separator problem, such that each component in $G$ has size at least $W+1$. If $|V|>2kW$ and $(G,k,W)$ is a yes-instance, then there exists a strictly reducible pair $(A,B)$ in $G$.
The fundamental structure utilized in the kernelization algorithm is again the BCD. The established relationships to COC are presented in the following lemma.
Lemma 4.44 ([DBLP:conf/esa/Casel0INZ21], Lemma 48 and 49).
Let $(G,k,W)$ be an instance of COC, and $(C,H,\mathfrak{R},f)$ a $W$-BCD in $G$. $(G,k)$ is a yes-instance if and only if $(G-(H\cup C),k-|A|)$ is a yes-instance. Furthermore, if $|\mathfrak{R}|+|H|>k$, then $(G,k)$ is a no-instance.
4.2 Kernelization Algorithm
In essence, the kernelization algorithm focuses on analyzing pairs $(A,B)$ in a $W$-BCD. We show that the head vertices $A$ have specific traits in a $W$-BCD, making it possible to locate them. Structurally, the algorithm operates as a bounded search tree, hence it is presented recursively. The input is an instance $(G,k,W)$ of COC, where $|V(G)|>2kW$ and each component of $G$ has at least $W+1$ vertices. The size limits for the input graph are not arbitrary; if $|V(G)|\leq 2kW$, there is nothing to do, and removing components smaller than $W+1$ is a safe reduction.
For $V^{\prime}\subseteq V$ we define $\texttt{sep}_{W}(V^{\prime})\in\mathds{N}$ as the cardinality of a minimum $W$-separator in $G[V^{\prime}]$, and $\texttt{arg\hskip 2.84544ptsep}_{W}(V^{\prime})\subset V$ as an argument suitable to this cardinality. For a graph $G^{\prime}\subseteq G$ let $G^{>W}(G^{\prime})$ be the graph obtained by removing all components of size at most $W$ from $G$. Lastly, for a $W$-BCD $(C,H,\mathfrak{R},f)$ we define :=$\mathfrak{R}^{\prime}:=\{R\in\mathfrak{R}\mid|R|>2W\text{ and }\texttt{sep}_{W}(R)=1\}$ and $S_{\mathfrak{R}^{\prime}}:=\bigcup_{R\in\mathfrak{R}^{\prime}}\texttt{arg\hskip 2.84544ptsep}_{W}(R)$ (uniqueness of $\texttt{arg\hskip 2.84544ptsep}_{W}(R)$ for $R\in\mathfrak{R}^{\prime}$ is shown in LABEL:lemma::UniqueSep).
Find reducible structures (
  1. 1.
    Compute a $W$-BCD $(C,H,\mathfrak{R},f)$ in $G$ and initialize $t=|H|+|\mathfrak{R}|$ and $S=\varnothing$.
  2. 2.
    Let $G^{\prime}=G^{>W}(G-S)$. If $G^{\prime}$ is an empty graph return a trivial yes-instance. Otherwise, compute a $W$-BCD $(C,H,\mathfrak{R},f)$ in $G^{\prime}$.
    1. (a)
      If $|H|+|\mathfrak{R}|>k$ return a trivial no-instance.
    2. (b)
      Let $\mathcal{Q}$ be the connected components of size at most $W$ in $G-S$. Let $A=S\cup H$ and $B=C\cup V(\mathcal{Q})$. Compute a DBE $(A_{1},A_{2},W,f,2W-1)$ in $G[A\cup B]$ using Theorem 1. If $A_{1}\neq\varnothing$, then terminate the algorithm and return $(G-(A_{1}\cup f^{-1}(A_{1})),k-|A_{1}|,W)$.
  3. 3.
    If the depth of the recursion is more than $\min(3t,k)$, then break.
  4. 4.
    For each $v\in H\cup S_{\mathfrak{R}^{\prime}}$:
    • Add $v$ to $S$ and recurse from step 2.
  5. 5.
    Return a trivial no-instance.
We split the correctness proof of Algorithm
Algorithm 3.
2kw into the following three lemmas:
Lemma 4.45.
Let $(G=(V,E),k,W)$ be a yes-instance and let $(A,B)$ be a minimal strictly reducible pair in $G$. Then, Algorithm
Lemma 4.46.
In the Algorithm
Lemma 4.47.
If Algorithm
If Lemma 4.45 holds, then the instance can be reduced by algorithm